### v1

# Mamo: a Mathematical Modeling Benchmark with Solvers

Abstract
Mathematical modeling involves representing real-world phenomena, systems, or problems using mathematical expressions and equations to analyze, understand, and predict their behavior. Given that this process typically requires experienced experts, there is an interest in exploring whether Large Language Models (LLMs) can undertake mathematical modeling to potentially decrease human labor. To evaluate of LLMs in mathematical modeling, we introduce a new benchmark, Mamo, that transcends traditional result-oriented assessments. Unlike conventional methods that primarily assess LLMs based on the accuracy of solutions to mathematical problems, our approach offers deeper insight into the modeling process itself. By focusing on the processes LLMs undertake rather than the correctness of their final solutions, Mamo pioneers a novel evaluation paradigm. This shift underscores the importance of understanding the inherent modeling capabilities of LLMs, paving the way for a more nuanced and comprehensive analysis of their problem-solving strategies. Our work marks a significant advancement in the field, suggesting a new direction for future research by emphasizing the evaluation of LLMs’ modeling processes over the mere correctness of answers. This benchmark not only facilitates a better understanding of LLMs’ mathematical modeling capabilities but also sets a new standard for evaluating their performance in complex problem-solving scenarios.
1 Introduction
Recent advancements in Large Language Models (LLMs) have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al. (2022); OpenAI (2023); Nijkamp et al. (2022); Tang et al. (2024). However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al. (2022). This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
Natural language captures human complexity and nuance, while mathematical language excels in precision and modeling the natural world (Giordano et al., 2013). Bridging their gap requires high-level cognitive skills akin to Artificial General Intelligence (AGI), a task at which LLMs show promise. However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem. They search through possible options to maximize or minimize an objective under given constraints if any. By using solves, The challenge becomes problem formulation under a given solver or a solver set. becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers. Recent studies have already combined LLMs with solvers Yang et al. (2023); Feng et al. (2023); Pan et al. (2023). In the field of optimization, AhmadiTeshnizi et al. (2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities.
We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of large language models (LLMs). Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies. We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
The contributions of this paper are as below:
  1. 1.
    Clarification of a Mathematical Modeling Framework for LLMs in Natural Language Domains: We have meticulously clarified the concept of mathematical modeling within Large Language Models (LLM), focusing on the realm of natural language challenges. This effort has culminated in the establishment of a comprehensive framework designed to guide the mathematical modeling process in LLMs, addressing both theoretical underpinnings and practical applications.
  2. 2.
    Innovation in Benchmarking Strategy through Solver Integration: Our research introduces a pioneering benchmarking strategy for the assessment of mathematical modeling process. By incorporating solvers into the benchmarking methodology, this strategy facilitates a rigorous evaluation process. This novel approach sets a new standard for the assessment in NLP research.
  3. 3.
    Creation of the Mamo Benchmark for Extensive Mathematical Modeling Assessment: We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions, focusing on areas such as ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks. With a total of 1059 meticulously curated questions, Mamo stands as a comprehensive tool for assessing the breadth and depth of mathematical modeling proficiency.
2 Related Work
Recent studies have made significant strides in applying LLMs to mathematical problem-solving, introducing innovative datasets for comprehensive evaluation. Frieder et al. (2024) and Yuan et al. (2023) have focused on datasets that test LLMs across various math problems and arithmetic expressions, respectively. Lewkowycz et al. (2022) explored training models using natural language paired with LaTeX-formatted mathematical content from arXiv, enhancing syntactic understanding. Zhang et al. (2024) developed the CARP dataset for computation-intensive challenges, while He et al. (2024) introduced OlympiadBench, a bilingual and multimodal benchmark for competition-level mathematical reasoning and proofs. Liu et al. (2024) constructed MathBench, which is designed with a carefully structured difficulty hierarchy and focuses on assessing the understanding of theoretical knowledge in LLMs. AhmadiTeshnizi et al. (2023a) constructed NLP4LP, a benchmark comprising 52 linear programming (LP) and mixed-integer linear programming (MILP) problems. These efforts collectively push the boundaries of LLMs in mathematical cognition, showcasing the growing complexity and specificity of tasks LLMs are being trained to tackle.
The pursuit of enhancing and devising novel solvers has seen a variety of research efforts, with a notable focus on the integration of LLMs. Yang et al. (2023) investigated the potential of employing LLMs as direct solvers, pioneering the exploration of LLMs’ autonomous problem-solving capabilities. He-Yueya et al. (2023) introduced a hybrid approach that marries an LLM with an external symbolic solver, specifically for equation solving, marking a significant advancement in combining neural and symbolic computation. Furthermore, Pan et al. (2023) developed LOGIC-LM, a cutting-edge framework that fuses LLMs with symbolic solvers, aimed at enhancing the solving of logical problems. In a closely related study, AhmadiTeshnizi et al. (2023a) introduced OptiMUS, an LLM-based agent explicitly crafted for formulating and solving optimization problems, thus demonstrating the broad applicability and potential of LLMs in tackling intricate computational challenges. These endeavors underscore the innovative directions being explored in the realm of solver development, leveraging the robust capabilities of LLMs.
3 Background
3.1 Mathematical Modeling
Giordano et al. (2013) emphasize the role of mathematical models as intermediaries that transform real-world problems into structured mathematical forms, facilitating the discovery of solutions. Despite LLMs’ proven competencies across various tasks, their capacity for developing mathematical models remains uncertain. This involves two key phases: model formulation, which demands a thorough comprehension of the problem, and model resolution, typically addressed by computational solvers. To accurately gauge LLMs’ potential in this domain, our strategy focuses on their ability to construct mathematical models, entrusting calculations to specialized solvers. This approach highlights LLMs’ creativity and generalization skills, simultaneously capitalizing on solvers’ computational efficiency.
Motivations to benchmark mathematical modeling for LLMs
We evaluate mathematical models for two reasons. The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can. Mathematical modeling is considered a highly advanced capability that we believe experts possess. This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence. On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al. (2021), etc., show that the performance has become saturated. Even a 7B model Yu et al. (2023b); Liu et al. (2023) (sometimes with a verifier-like model Yu et al. (2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4. Therefore, we believe there is a need to construct more challenging mathematical datasets. Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
3.2 Using Solvers in Mathematical Modeling
The model generated by the LLM is then subjected to verification against real-world data to assess its validity, ensuring it addresses the problem logically and sensibly. The solver, an algorithmic tool designed to find solutions to mathematical models, is also involved in this process and it is also beneficial for verification.
Different types of solvers are specialized for various classes of problems, such as optimization solvers. However, they share the common property: They are all algorithmic tools designed to find solutions to mathematical models.
Rationale of Using solver in Mathematical Modeling for LLMs These solvers operationalize the abstract constructs created by LLMs, translating them into tangible outputs that can be analyzed and validated against real-world data and phenomena. The solver is a crucial instrument that resolves the model presented by the LLM, and by comparing the solution it provides with the correct answer, we assess the accuracy of the mathematical model generated by the LLM.
4 Philosophy of Mamo Benchmark
4.1 Methodology: exact answer verification with solvers
We utilize solver to resolve this mathematical model. The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps. By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$). This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
Figure 1: The pipeline to use exact answer verification via an additional solver.
4.2 Scope of Mathematical Modeling
We construct the category of Mathematical Modeling by referring to Giordano et al. (2013). In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Tab. 1.
Category In Mamo
Changes and Differences
Proportionality and Geometric Similarity
Optimization Problem Modeling
Differential Equations
Probabilistic Modeling
Table 1: The taxonomy of mathematical modeling. In Mamo benchmark, we only include the categories where sophisticated Solvers is available.
Given the broad spectrum of mathematical modeling and the availability of specific solvers, our benchmark is meticulously designed to encompass domains where these tools can be effectively utilized. This section outlines the rationale for our focus on certain areas of mathematical modeling, based on solver availability and the clarity of the modeling process.
Availability of Sophisticated Solvers: Advanced optimization solvers, such as COPT Ge et al. (2023), along with Python libraries for solving ordinary differential equations (ODEs), offer a strong foundation for testing the capabilities of Large Language Models (LLMs). These tools enable a detailed and automated evaluation of the LLMs’ abilities to abstract, formulate, and accurately solve mathematical problems within these specific domains.
Challenges in Modeling Changes and Differences, Proportionality and Geometric Similarity: Although simpler tools like calculators could address problems related to modeling changes and differences, proportionality and geometric similarity, the lack of a defined, structured modeling process complicates the implementation of our final-state-approach benchmark. Our objective is to assess the LLMs’ skill in developing comprehensive mathematical models that adhere to a solvable framework. The absence of a distinct modeling procedure in these areas impedes a standardized and objective assessment of an LLM’s modeling efficiency.
Following the considerations mentioned above, we have developed our benchmark in the fields of optimization, and ODE, due to the lack of advanced and universal Partial Differential Equations (PDE) solvers. In the optimization segment, our focus is on Linear Programming (LP) and Mixed-Integer Linear Programming (MILP) problems. Compared to NLP4LP, introduced by AhmadiTeshnizi et al. (2023b) and focusing exclusively on optimization, our benchmark includes a larger number of problems. This benchmark set is highly scalable, with potential future expansions into nonlinear optimization and probabilistic modeling. Additionally, as new and advanced solvers emerge, it could be expanded into areas like Partial Differential Equations (PDE) and beyond.
5 Data Collection and Synthesis
5.1 Data Selection and Source Credibility
The data set for our benchmark consists of a blend of manually selected and GPT-generated questions. All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance.
Annotator Qualifications
The benchmarking process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines.  [^2] See the details of the annotator qualifications in Appendix. B:
5.1.1 Manually Selected Questions
  • Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010). The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario. The source ensure the correctness of the answer of each question.
  • Rewriting Process: Each question is meticulously rewritten to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology. This process also circumvents potential licensing issues, as the original content is significantly transformed.
Data Synthesis
The process of annotating data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems. This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations. See the example in Appendix. E.
Cross Validation
GPT-4 is employed to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach, ensuring the benchmark’s focus remains on the LLM’s modeling capabilities. Moreover, we employ a comprehensive review process, re-examining each question for consistency and correctness, thereby ensuring the quality of our dataset. See the review details in Appendix. A.
5.2 Data Statistics and Visualization
In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
See the word cloud of the combined data in Figure 2
Figure 2: The word cloud of the data after filtering the requirement words such as ’significant figures’, ’rounded’
5.3 Guidance of Data Synthesis
The synthesis of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs). The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification. The following are the key considerations in our benchmark guidance:
Property 1.
Final-State Approach Solvability: Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
Property 2.
Numerical Answers: The answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
Property 3.
Significant Figures and Precision: Each question will explicitly state the required significant figures or the level of numerical precision for the answer. This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
Property 4.
Modeling Capacity Test: To specifically test the LLM’s modeling capabilities:
  • For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$.
  • For optimization problems, questions should aim to find the optimal value.
These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem.
Property 5.
Real-World Problem Context: Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
6 Evaluation
6.1 Evaluation Protocol
6.1.1 Evaluation Details
Our evaluation protocol is meticulously designed to test the LLMs’ mathematical modeling process. For ODE problems, the LLM is prompted to generate Python code from a natural language description. The code is executed, and its output is compared with the correct answer to assess accuracy. In optimization problems, the LLM is asked to express the model in .lp format, after which the COPT solver is used to find the optimal value for comparison with the correct answer.
Model output
In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy). In the optimization sections, we prompt LLMs to generate the standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable. Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities. To better adhere to the format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix. (see Appendix C and Appendix. G).
Mitigation of Formatting Errors
It is essential to distinguish errors in modeling from those in coding or file formatting. Therefore, if the LLM’s output contains syntax errors or incorrect .lp formatting, we need to rectify these issues without altering the underlying logic of the code or model. This step ensures that our evaluation focuses solely on the modeling process. To achieve this, we use a language model to fix any syntax errors, referring to it as the code modifier. In our experiments, we use the original language model and GPT-4-0613 as code modifiers. For the sake of reproducibility, we choose the orignal LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models. See the comparison of the result from different code modifier in Section 6.3 and the prompts in Appendix.  G. Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like " “‘python “‘ " that will affect our testing.
6.1.2 Evaluation Metrics
Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies. Let $O$ denote the output from the LLM, and $A$ denote the standard answer. The comparison is defined by the following criteria  333 See the evaluation script in Appendix. F
Precision Adjustment for Floating-Point Numbers If the standard answer, $A$, is a floating-point number, the precision adjustment involves calculating the number of significant digits, $n$, after the decimal point. Both $O$ and $A$ are then scaled by $10^{n}$ for comparison.
General Comparison Criterion The overall correctness is determined based on the following conditions, applicable to both adjusted and original values:
$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{ or }|O-A|\leq 1,\\ 0&\text{otherwise.}\end{cases}$ (1)
This structured evaluation ensures that the LLM’s output is accurately assessed against the standard answer with consideration for the specific requirements of numerical precision and the inherent variability in computational answers.
6.2 Benchmarking Results
Models ODE Optimization
First order Second order System All Easy LP Complex LP
GPT-4o 69.39% 73.64% 42.50% 67.63% 87.27% 23.22%
GPT-4o 69.90% 74.55% 42.50% 68.21% 87.42% 32.23%
GPT-4-turbo 66.84% 69.09% 42.50% 65.12% 87.88% 23.70%
GPT-4-turbo 67.86% 70.00% 42.50% 65.99% 87.88% 30.33%
GPT-4 60.71% 65.45% 42.50% 60.12% 86.50% 22.27%
GPT-4 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
GPT-3.5-turbo 17.35% 16.36% 7.50% 15.90% 81.29% 9.95%
GPT-3.5-turbo 25.51% 31.82% 15.00% 26.30% 84.82% 9.95%
Claude-3-Sonnet 53.57% 61.82% 30.00% 53.47% 83.59% 25.12%
Claude-3-Sonnet 55.61% 63.64% 35.00% 55.78% 84.20% 25.59%
Claude-3-Haiku 42.35% 60.91% 20.00% 45.66% 78.53% 19.91%
Claude-3-Haiku 52.04% 63.64% 22.50% 52.31% 85.89% 21.33%
Gemini-1-pro 26.53% 35.45% 2.50% 26.59% 78.99% 15.31%
Gemini-1-pro 35.71% 39.09% 7.50% 33.53% 80.06% 15.31%
Gemini-1.5-pro 66.33% 66.36% 47.50% 64.16% 85.28% 30.81%
Gemini-1.5-pro 67.35% 66.36% 52.50% 65.32% 85.28% -
DeepSeek-v2 40.82% 45.45% 32.50% 41.33 % 86.50% 6.16%
DeepSeek-v2 42.35% 64.55% 32.50% 48.27% 86.66% 6.16%
DeepSeek-coder 32.14% 49.09% 15.00% 35.55% 75.77% 18.96%
DeepSeek-coder 34.18% 49.09% 15.00% 36.71% 76.99% 18.96%
Table 2: Evaluation result on Mamo Benchmark. The remark $\dagger$ indicates it additionally uses the original LLM to rectify syntax error of the code and .lp file generated by the corresponding LLM. Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09 and GPT-4o denotes GPT-4o-2024-05-13, while GPT-4 refers to GPT-4-0613
Our evaluation includes both closed-source LLMs and open-source LLMs. The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^4] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^5] https://deepmind.google/technologies/gemini/. The open-source LLMs include DeepSeek-v2 DeepSeek-AI (2024) and DeepSeek-coder Guo et al. (2024). Table 2 presents the evaluation results of different language models on the Mamo Benchmark. The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (indicated by the dagger symbol ). Table 3 displays the performance results when GPT-4-0613 is selected as the code modifier.
Models ODE Optimization
First order Second order System All Easy LP Complex LP
GPT-4o †† 69.90% 74.55% 42.50% 68.21% 87.42% 29.86%
GPT-4-turbo †† 67.85% 70.00% 42.50% 65.99% 88.19% 28.44%
GPT-4 †† 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
GPT-3.5-turbo †† 47.96% 60.91% 20.00% 48.84% 85.43% 10.43%
Claude-3-Sonnet †† 58.16% 66.36% 35.00% 59.09% 85.12% 28.44%
Claude-3-Haiku †† 53.06% 65.45% 25.00% 53.76% 86.81% 21.33%
Gemini-1-pro †† 43.37% 54.55% 15.00% 43.64% 80.37% 16.27%
Gemini-1.5-pro †† 67.35 % 67.27% 52.50% 65.61% 85.28% 34.12%
DeepSeek-v2 †† 56.12% 68.18% 32.50% 57.23% 86.81% 6.16%
DeepSeek-coder †† 33.67% 49.09% 17.50% 36.71% 78.83% 18.96%
[…]
The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks. Specifically, GPT-4o achieves the highest performance in ODE tasks with 69.90% in first-order and 74.55% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 32.23% in complex LP scenarios when using the LLM to rectify syntax errors. Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with 48.27% overall in ODE tasks and 86.66% in easy LP tasks when enhanced with syntax correction, highlighting its potential despite being open-source.
The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models. For instance, GPT-3.5 Turbo’s performance in ODE tasks increased from 15.90% to 26.30% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 45.66% to 52.31%. Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 41.33% to 48.27% when enhanced with syntax correction.
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 48.27% with self-improvement to 57.23% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
6.3 On the format errors and code modifiers
Models ODE Optimization
format raw self-improved GPT-4-improved raw self-improved GPT-4-improved
GPT-4o 96.82% 99.42% 99.42% 92.93% 97.91% 97.91%
GPT-4-turbo 96.82% 99.13% 98.84% 93.86% 97.80% 97.56%
GPT-4 94.51% 97.69% 97.69% 90.96% 95.13% 95.13%
GPT-3.5-turbo 32.37% 60.40% 94.22% 87.95% 89.80% 96.76%
Claude-3-Sonnet 84.68% 93.06% 97.40% 91.89% 93.97% 96.76%
Claude-3-Haiku 78.03% 92.77% 98.27% 82.61% 92.58% 96.29%
Gemini-1.0-pro 56.36% 69.94% 88.73% 88.06% 89.11% 94.21%
Gemini-1.5-pro 93.64% 98.27% 98.84% 95.48% - 96.87%
deepseek-v2 70.81% 83.81% 96.24% 90.27% 92.00% 94.90%
deepseek-coder 88.44% 93.64% 95.95% 92.58% 96.29% 98.84%
Table 4: Correct format rate on raw, self-improved, and GPT4-improved format.
The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different improvement methods. The table considers three scenarios: no improvement, self-improvement, and improvement provided by GPT-4.
The GPT-3.5-turbo model shows the most dramatic improvement when enhanced by GPT-4, jumping from a 32.37% success rate to 94.22%. Interestingly, Claude-3 models (Sonnet, and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4, highlighting the compatibility of GPT-4’s improvement methods with other models. In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
The improvement from GPT-4 is generally superior to both self-improvement and no improvement. This indicates that GPT-4 can serve as a more effective and fair code modifier. By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors. Notably, in ODE tasks, the correct format rate for all models, except one, is generally greater than 94%, with the lowest being 88.73%. In Optimization tasks, the correct format rate is even higher, with the lowest rate being 94.21%.
7 Conclusion and Future Directions
In this article, we introduce Mamo, a novel benchmark that utilizes solvers to evaluate the mathematical modeling capabilities of large language models (LLMs). We propose a shift in the evaluation paradigm from traditional outcome-oriented evaluation to process-oriented evaluation. This shift allows us to scrutinize the modeling process undertaken by LLMs, providing a more detailed and comprehensive assessment of their mathematical problem-solving abilities. Our evaluation focuses more on the mathematical modeling capabilities of the models, leaving the process of solving abstract mathematical problems to the solvers, thereby avoiding the impact of errors such as computational capabilities of the models on their actual modeling abilities.
Our benchmark represents an important step in understanding the fundamental modeling capabilities of LLMs. By focusing on the modeling process rather than the correctness of the final solution, we are able to gain valuable insights into the strengths and limitations of current LLMs. This approach provides new directions for future research, emphasizing the importance of the modeling process in solving mathematical problems.
The introduction of Mamo paves new avenues for future research. The benchmark can be expanded to include a wider variety of mathematical problems and solvers, broadening its applicability and providing a more comprehensive evaluation of the mathematical modeling capabilities of LLMs. Insights gained from the benchmark can be used to guide the development of future LLMs. By understanding the strengths and weaknesses of current models in solving mathematical problems, researchers can design new models that better handle these tasks. We will extend our model list to other models (Llama2 series, Mistral series etc) in our further study.
Limitations
While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research. The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning. This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities. Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
Ethics Statement:
The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes. Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups. This includes but is not limited to issues of privacy, security, and the potential for misuse of the model. Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.
Appendix A Cross Review
In ODE part, we have the following review process:
  1. 1.
    We employed GPT-4 to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach. There were only 10 out of 383 questions are invalid. All were filtered, 373 questions were left.
  2. 2.
    Then we conducted a thorough evaluation of the questions. Our dataset initially comprised $373$ questions. We identified $29$ questions that lacked sufficient information, $21$ that had incorrect answers, and $18$ that featured unclear statements. Following our assessment, $27$ questions were deleted, and $41$ were corrected to ensure clarity and correctness.
Finally, we have 346 problems in ODE part.
We have also conducted the cross-review process. See the details in Appendix.  A.1 and the results on Appendix.  A.1.1.
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, 8 of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found synthetized by wrong mathematical models. All of them were corrected.
A.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section B) were tasked with assessing 50 questions each. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 6.1.2. Here are our results:
A.1.1 Results
The effectiveness of the review process was quantified using the metrics defined in Subsection 6.1.2, focusing on the accuracy and inter-rater reliability among reviewers. The statistical outcomes are summarized as follows:
  • Average Cohen’s Kappa: The average Cohen’s Kappa across all reviewer pairs was 0.60, indicating a moderate to substantial agreement. This suggests that reviewers generally agreed on the classification of answers, though there were variations in some cases.
  • Minimum and Maximum Cohen’s Kappa: The minimum Kappa value recorded was 0.50, and the maximum was 0.71. The spread in these values highlights areas where alignment and training could potentially enhance consistency.
  • Accuracy Rates: The individual accuracy rates were as follows:
    • Reviewer 1: 90.0%
    • Reviewer 2: 84.0%
    • Reviewer 3: 88.0%
    • Reviewer 4: 74.0%
    These rates reflect the precision with which each reviewer matched the standard answers, including their recognition of problematic questions correctly labeled as “error.”
These results provide insights into the overall effectiveness and areas of improvement for the cross-review process, particularly in terms of aligning understanding and interpretation of the evaluation criteria among reviewers.
Appendix B Annotator Qualifications
The benchmarking process benefits from the expertise of reviewers with robust backgrounds in mathematics, underscored by their academic accomplishments. The team is composed of individuals whose education encompasses critical mathematical disciplines, including calculus (I and II), linear algebra (both introductory and advanced levels), optimization, probability, and ordinary differential equations. The average Grade Point Average (GPA) across these foundational courses is approximately 3.89 out of a 4.0 scale. This high level of academic achievement demonstrates the reviewers’ strong grasp of essential mathematical concepts and their application. Furthermore, the team is enriched by members who hold Ph.D. degrees in mathematics, further solidifying the depth of expertise and analytical skills available for the benchmarking process.
B.1 Qualification of Reviewers
We select the reviewers mentioned in Appendix  A.1 based on the following criteria:
  1. 1.
    Undergraduate students who have taken fundamental mathematical courses (Calculus and Linear Algebra) and have an average GPA of at least 3.5/4.0.
  2. 2.
    Have completed or are currently taking courses in Ordinary Differential Equations (ODE).
  3. 3.
    Meet one of the following English proficiency requirements: TOEFL > 80, IELTS > 6.5, a minimum grade of B+ in all English classes, or a Gaokao English score > 120.
These criteria ensure their ability to read questions in English and solve ODE problems.
Appendix C Few-shot Experiment
To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance. In our analysis, we tested the performances of 11 models on ODE problems using prompts with 0 shot, 1 shot, 3 shots, 5 shots, and 10 shots. The results, shown in Figure  3, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots. Similarly, we performed sensitivity tests on optimization problems, assessing the accuracy of 10 models with prompts of 0 shot, 1 shot, 3 shots, and 5 shots. The results are depicted in Figure  4.
Figure 3: Results of models accuracy and shots on ODE problems
Figure 4: Results of models accuracy and shots on optimization problems
Additionally, we specifically tested the 0-shot performance of GPT-4-0613 on all questions in Easy LP and Complex LP categories. The accuracy of GPT-4-0613 was 66.56% in Easy LP and 14.69% in Complex LP, resulting in an overall LP accuracy of 53.65%, consistent with the results obtained from sampling.
Appendix D Testing Process
We use 3-shot learnging in testing, see examples in  G
Figure 5: Example of testing process in optimization
Figure 6: Example of testing process in ODE
Appendix E Data Synthesis
Figure  7 shows the example of synthesis process in ODE.
Figure 7: Example of synthesis process in ODE
Appendix F Evaluation script
def comp(output, standard_answer):
    dif = abs(float(output) - float(standard_answer))
    if float(standard_answer) == 0:
        rate = dif * 1e-4
    else:
        rate = dif / float(standard_answer)
    if abs(rate) <= 1e-4:
        return 1
    else:
        return 0
def compare_output_with_standard(output, standard_answer):
    try:
        float_output = float(output)
    except ValueError:
        return False
    if ’.’ in standard_answer:
        digit = len(standard_answer.split(’.’)[1])
        s_ans = float(standard_answer) * 10 ** digit
        ans = float_output * 10 ** digit
        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
    else:
        s_ans = float(standard_answer)
        ans = float_output
        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
Appendix G Prompts
Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: Example_1: (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
Figure 8: The prompt for testing in optimization part. [question] refers to a question in the benchmark. […] refers to the examples. We use 3-shot in testing
Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: ```$python$ def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 `````````` ` ` 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the examples: […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: ′′′ Here’s the Python code that solves the given problem and meets the specified requirements: `````````` ` `python `````````` ` ` This code uses the `$solve_{i}vp`$ function from SciPy to solve the initial value problem for the given differential equation. The `$round_{t}o_{s}ignificant_{f}igures`$ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. ′′′ Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
Figure 9: The prompt for testing in ODE part. [question] refers to a question in the benchmark. […] refers to the examples.
You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the code. Your response should be entirely in Python code, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No other comments outside the code are allowed. The followings are forbidden: ”’ The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘. The round function is missing a closing parenthesis. Here is the corrected version: “‘python “‘ ”’ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the code: [code] The correct code is:
Figure 10: The prompt for fixing syntax error in ODE. [error] refers to the message when executing the Python code. While [code] refer to the code which contains syntax error
You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the lp file. Your response should be entirely in .lp format, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No comments are allowed. The followings are forbidden: ”’ This is the correct .lp file: “‘lp “‘ ”’ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the lp: [lp_code] Generate the correct .lp format, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality. For example, make a + b + c <= d into a + b + c - d <= 0. Ensure there is a space between variables and their coefficients (coefficients should be numerical). The correct lp code is:
Figure 11: The prompt for fixing syntax error in LP. [lp_code] refer to the lp code which contains syntax error

### v2

@@
 1 Introduction
-Recent advancements in Large Language Models (LLMs) have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al.
+Recent advancements in Large Language Models have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al.
 (2022); OpenAI (2023); Nijkamp et al.
@@
 They search through possible options to maximize or minimize an objective under given constraints if any.
-By using solves, The challenge becomes problem formulation under a given solver or a solver set.
+By using solvers, The challenge becomes problem formulation under a given solver or a solver set.
 becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers.
@@
 This implies the potential for testing LLM’s modeling capabilities.
-We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of large language models (LLMs).
+We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of LLMs.
 Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies.
 We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
-The contributions of this paper are as below:
-  1. 1.
-    Clarification of a Mathematical Modeling Framework for LLMs in Natural Language Domains: We have meticulously clarified the concept of mathematical modeling within Large Language Models (LLM), focusing on the realm of natural language challenges.
-This effort has culminated in the establishment of a comprehensive framework designed to guide the mathematical modeling process in LLMs, addressing both theoretical underpinnings and practical applications.
-  2. 2.
-    Innovation in Benchmarking Strategy through Solver Integration: Our research introduces a pioneering benchmarking strategy for the assessment of mathematical modeling process.
-By incorporating solvers into the benchmarking methodology, this strategy facilitates a rigorous evaluation process.
-This novel approach sets a new standard for the assessment in NLP research.
-  3. 3.
-    Creation of the Mamo Benchmark for Extensive Mathematical Modeling Assessment: We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities.
-Mamo encompasses a wide array of modeling questions, focusing on areas such as ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks.
-With a total of 1059 meticulously curated questions, Mamo stands as a comprehensive tool for assessing the breadth and depth of mathematical modeling proficiency.
+The contributions of this paper are as below: (1) We introduce a new benchmarking strategy to assess mathematical modeling through Solvers, facilitating a rigorous evaluation process.
+(2) We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities.
+Mamo encompasses a wide array of modeling questions (With a total of 1209 meticulously curated questions), including ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks.
 2 Related Work
@@
 This approach highlights LLMs’ creativity and generalization skills, simultaneously capitalizing on solvers’ computational efficiency.
-Motivations to benchmark mathematical modeling for LLMs
-We evaluate mathematical models for two reasons.
+Motivationsto Benchmark Mathematical Modeling for LLMs   We evaluate mathematical models for two reasons.
 The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can.
@@
 The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps.
-By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$).
+By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$), since typically the wrong ${\widehat{M}}$ will result in wrong ($\widehat{A}$), like the examples in Appendix  G.
 This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
@@
 (2013).
-In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Tab.
-1.
-Category In Mamo
-Changes and Differences
-Proportionality and Geometric Similarity
-Optimization Problem Modeling
-Differential Equations
-Probabilistic Modeling
-Table 1: The taxonomy of mathematical modeling. In Mamo benchmark, we only include the categories where sophisticated Solvers is available.
+In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Table 7 in Appendix F.
 Given the broad spectrum of mathematical modeling and the availability of specific solvers, our benchmark is meticulously designed to encompass domains where these tools can be effectively utilized.
@@
 Additionally, as new and advanced solvers emerge, it could be expanded into areas like Partial Differential Equations (PDE) and beyond.
-5 Data Collection and Synthesis
+5 Data Collection
 5.1 Data Selection and Source Credibility
-The data set for our benchmark consists of a blend of manually selected and GPT-generated questions.
+The dataset for our benchmark consists of a blend of manually selected and GPT-synthetized questions.
 All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance.
-Annotator Qualifications
-The benchmarking process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines.
-[^2] See the details of the annotator qualifications in Appendix.
-B:
-5.1.1 Manually Selected Questions
-  • Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010).
+The data collection and synthesis process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines  [^2] See the details of the annotator qualifications in Appendix C.
+Manually Selected Questions
+The manually selected questions are high quality and reliable.
+(1)Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010).
 The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario.
 The source ensure the correctness of the answer of each question.
-  • Rewriting Process: Each question is meticulously rewritten to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology.
-This process also circumvents potential licensing issues, as the original content is significantly transformed.
-Data Synthesis
-The process of annotating data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity.
+(2) Transformation: Each question is meticulously transformed to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology.
+Question Formulation and Verification
+The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity.
 This is followed by the computation of answers to establish a ground truth.
@@
 This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations.
-See the example in Appendix.
-E.
-Cross Validation
-GPT-4 is employed to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach, ensuring the benchmark’s focus remains on the LLM’s modeling capabilities.
-Moreover, we employ a comprehensive review process, re-examining each question for consistency and correctness, thereby ensuring the quality of our dataset.
-See the review details in Appendix.
-A.
-5.2 Data Statistics and Visualization
-In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
-The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
-See the word cloud of the combined data in Figure 2
-Figure 2: The word cloud of the data after filtering the requirement words such as ’significant figures’, ’rounded’
-5.3 Guidance of Data Synthesis
-The synthesis of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs).
+See the example in Appendix K.
+5.2 Data Quality Checking
+In the ODE part, we employed GPT-4 to confirm the solvability of newly generated questions, filtering out invalid ones.
+A thorough evaluation identified questions lacking information, with incorrect answers, and unclear statements.
+Consequently, problematic questions were deleted or corrected, resulting in a refined dataset.
+In the optimization section, particularly with the easy LP part, invalid questions were filtered out, and those with incorrect answers or misleading descriptions were corrected.
+In the complex LP part, questions formulated with wrong mathematical models were corrected.
+The summary of the review process is presented in Table 1.
+Category Initial Questions Filtered Final Questions
+ODE Part 383 37 346
+Easy LP Part 688 48 640
+Complex LP Part 211 0 211
+Table 1: Summary of the review process.
+Reviewer Accuracy Rate
+Reviewer 1 90.0%
+Reviewer 2 84.0%
+Reviewer 3 88.0%
+Reviewer 4 74.0%
+Table 2: Accuracy of reviewers.
+For the cross-review process, 50 questions from the ODE dataset were selected for deeper analysis by four independent reviewers.
+Reviewers evaluated the questions, marking responses as either numeric answers or "error" if the question was problematic.
+The effectiveness of this review was quantified using metrics like Cohen’s Kappa to measure inter-rater reliability and accuracy rates for individual reviewers.
+The average Cohen’s Kappa was 0.60, indicating moderate to substantial agreement among reviewers, with individual Kappa values ranging from 0.50 to 0.71.
+Accuracy rates are moderately high as shown in Tables 2.
+These results highlight both the effectiveness of the review process and areas where alignment and training can improve consistency and accuracy among reviewers.
+See the review details in Appendix B.
+5.3 Guidance of Data Collection
+The collection of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs).
 The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification.
@@
 Property 1.
-Final-State Approach Solvability: Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
+Final-State Approach Solvability  Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
 Property 2.
-Numerical Answers: The answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
-Property 3.
-Significant Figures and Precision: Each question will explicitly state the required significant figures or the level of numerical precision for the answer.
-This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
-Property 4.
-Modeling Capacity Test: To specifically test the LLM’s modeling capabilities:
-  • For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$.
+Unified and Numerical Answers: To handle potentially various valid solutions for one problem: For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$.
 For optimization problems, questions should aim to find the optimal value.
 These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem.
-Property 5.
-Real-World Problem Context: Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
+Besides, the answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
+Property 3.
+Significant Figures and Precision  Each question will explicitly state the required significant figures or the level of numerical precision for the answer.
+This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
+To assess the LLM’s capability to abstract mathematical models from real-world scenarios, we introduce Property 4.
+Property 4.
+Real-World Problem Context  Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
+5.4 Data Statistics
+In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
+The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
+See the word cloud of the combined data in the Appendix L.
 6 Evaluation
@@
 To better adhere to the format, we conducted tests using 3-shot learning.
-Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix.
-(see Appendix C and Appendix.
-G).
+Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix (see Appendix I and Section O).
+Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like ‘ “‘ python “‘ ’ that will affect our testing.
 Mitigation of Formatting Errors
-It is essential to distinguish errors in modeling from those in coding or file formatting.
-Therefore, if the LLM’s output contains syntax errors or incorrect .lp formatting, we need to rectify these issues without altering the underlying logic of the code or model.
+It is essential to distinguish errors in modeling from those in coding or file formatting [^3] Appendix  H displays the example of formating errors instead of modeling errors..
+Therefore, if the LLM’s output contains syntax errors in python code or .lp code, we need to rectify these issues without altering the underlying logic of the code.
 This step ensures that our evaluation focuses solely on the modeling process.
-To achieve this, we use a language model to fix any syntax errors, referring to it as the code modifier.
+To achieve this, we use a language model to fix any syntax error, referring to it as the code modifier.
 In our experiments, we use the original language model and GPT-4-0613 as code modifiers.
-For the sake of reproducibility, we choose the orignal LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models.
-See the comparison of the result from different code modifier in Section 6.3 and the prompts in Appendix.
-G.
-Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like " “‘python “‘ " that will affect our testing.
+For the sake of reproducibility, we choose the original LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models.
+See the comparison of the result from different code modifiers in Section 6.3, and prompts in Appendix  O.
 6.1.2 Evaluation Metrics
@@
 Let $O$ denote the output from the LLM, and $A$ denote the standard answer.
-The comparison is defined by the following criteria  333 See the evaluation script in Appendix.
-F
-Precision Adjustment for Floating-Point Numbers If the standard answer, $A$, is a floating-point number, the precision adjustment involves calculating the number of significant digits, $n$, after the decimal point.
-Both $O$ and $A$ are then scaled by $10^{n}$ for comparison.
-General Comparison Criterion The overall correctness is determined based on the following conditions, applicable to both adjusted and original values:
-$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{ or }|O-A|\leq 1,\\ 0&\text{otherwise.}\end{cases}$ (1)
-This structured evaluation ensures that the LLM’s output is accurately assessed against the standard answer with consideration for the specific requirements of numerical precision and the inherent variability in computational answers.
+The comparison is defined by the following criteria  444 See the evaluation script in Appendix M
+General Comparison Criterion Let $n$ denotes the decimal places of $A$ ($n=0$ if $A$ is integer). The Answer $O$ is correct if
+$\left|\frac{O-A}{A}\right|<10^{-4}\text{ or }|O-A|<\min\{10^{-n},10^{-2}\}$ (1)
+This metric is designed [^5] See the case study in Appendix D to account for the precision errors inherent in solvers and the specified significant figures in questions, ensuring that the LLM’s output is accurately evaluated against the standard answer with allowances for numerical precision and variability in computational responses.
 6.2 Benchmarking Results
-Models ODE Optimization
-First order Second order System All Easy LP Complex LP
-GPT-4o 69.39% 73.64% 42.50% 67.63% 87.27% 23.22%
-GPT-4o 69.90% 74.55% 42.50% 68.21% 87.42% 32.23%
-GPT-4-turbo 66.84% 69.09% 42.50% 65.12% 87.88% 23.70%
-GPT-4-turbo 67.86% 70.00% 42.50% 65.99% 87.88% 30.33%
-GPT-4 60.71% 65.45% 42.50% 60.12% 86.50% 22.27%
-GPT-4 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
-GPT-3.5-turbo 17.35% 16.36% 7.50% 15.90% 81.29% 9.95%
-GPT-3.5-turbo 25.51% 31.82% 15.00% 26.30% 84.82% 9.95%
-Claude-3-Sonnet 53.57% 61.82% 30.00% 53.47% 83.59% 25.12%
-Claude-3-Sonnet 55.61% 63.64% 35.00% 55.78% 84.20% 25.59%
-Claude-3-Haiku 42.35% 60.91% 20.00% 45.66% 78.53% 19.91%
-Claude-3-Haiku 52.04% 63.64% 22.50% 52.31% 85.89% 21.33%
-Gemini-1-pro 26.53% 35.45% 2.50% 26.59% 78.99% 15.31%
-Gemini-1-pro 35.71% 39.09% 7.50% 33.53% 80.06% 15.31%
-Gemini-1.5-pro 66.33% 66.36% 47.50% 64.16% 85.28% 30.81%
-Gemini-1.5-pro 67.35% 66.36% 52.50% 65.32% 85.28% -
-DeepSeek-v2 40.82% 45.45% 32.50% 41.33 % 86.50% 6.16%
-DeepSeek-v2 42.35% 64.55% 32.50% 48.27% 86.66% 6.16%
-DeepSeek-coder 32.14% 49.09% 15.00% 35.55% 75.77% 18.96%
-DeepSeek-coder 34.18% 49.09% 15.00% 36.71% 76.99% 18.96%
-Table 2: Evaluation result on Mamo Benchmark.
-The remark $\dagger$ indicates it additionally uses the original LLM to rectify syntax error of the code and .lp file generated by the corresponding LLM.
-Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09 and GPT-4o denotes GPT-4o-2024-05-13, while GPT-4 refers to GPT-4-0613
-Our evaluation includes both closed-source LLMs and open-source LLMs.
+Table 3: Evaluation Results on the Mamo Benchmark.
+The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 6.1.1).
+The "Overall" score represents the weighted average of correct rates, weighted by question count.
+The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier.
+The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
+Models ODE Optimization Overall (%)
+First Second System Easy Complex
+order (%) order (%) (%) LP (%) LP (%)
+GPT-4o 67.86 +0.51 36.36 +0.91 40.00
++0.00
+87.27
++0.15
+22.75
++8.53
+66.67 +1.73
+GPT-4-turbo 64.80
++1.02
+32.73
++0.91
+40.00
++0.00
+87.88 +0.00 23.22
++6.64
+66.34
++1.41
+GPT-4 59.18
++1.02
+28.18
++0.91
+40.00
++2.50
+86.50
++0.46
+21.08
++3.09
+63.81
++1.12
+GPT-3.5-turbo 16.33
++6.12
+7.27
++4.55
+10.00
++5.00
+81.29
++3.53
+9.48
++0.00
+49.13
++3.48
+Claude-3-Opus 55.61
++3.57
+33.64
++1.81
+45.00
++2.50
+83.74
++0.31
+9.48
++0.47
+60.38
++1.08
+Claude-3-Sonnet 52.04
++2.04
+24.55
++0.90
+27.50
++5.00
+83.59
++0.61
+24.64
++0.48
+60.96
++0.99
+Claude-3-Haiku 41.33
++9.18
+23.64
++0.91
+17.50
++2.50
+78.53
++7.36
+17.54
++1.42
+54.84
++5.87
+Gemini-1-pro 25.00
++8.16
+13.64
++0.00
+2.50
++2.50
+78.83
++1.08
+14.69
++0.00
+50.45
++1.99
+Gemini-1.5-pro 64.80
++1.02
+30.00
++0.00
+45.00 +5.00 85.28
++0.00
+30.81 +5.21 66.09
++1.24
+DeepSeek-v2 39.29
++1.02
+20.00
++9.09
+30.00
++0.00
+86.50
++0.16
+3.79
++0.00
+56.49
++1.08
+DeepSeek-coder-33B 28.57
++2.04
+11.82
++0.00
+12.50
++0.00
+75.77
++1.22
+18.96
++0.00
+50.29
++0.99
+DeepSeek-math-7b-base 3.57
++0.00
+2.73
++0.00
+2.50
++0.00
+49.39
++1.22
+6.64
++0.00
+28.70
++0.66
+DeepSeek-math-7b-rl 1.02
++0.51
+1.82
++0.00
+5.00
++0.00
+20.55
++0.62
+0.00
++0.47
+11.58
++0.50
+Llama-3-8b 0.51
++0.51
+0.00
++0.00
+0.00
++0.00
+55.37
++0.46
+6.64
++0.00
+31.10
++0.33
+Llama-3-70b 4.08
++0.51
+1.82
++1.82
+2.50
++0.00
+66.10
++0.62
+9.00
++0.00
+38.13
++0.58
+Qwen-1.5-72b-Chat 5.61
++6.12
+4.55
++0.00
+5.00
++7.50
+65.03
++6.29
+5.69
++0.00
+37.55
++4.64
+Qwen-1.5-110b-Chat 28.57
++4.08
+6.36
++1.82
+7.50
++0.00
+65.18
++10.89
+13.27
++2.37
+43.93
++6.11
+Mixtral-8x7b-instruct 23.47
++4.59
+4.55
++1.81
+10.00
++0.00
+67.79
++2.15
+9.95
++0.00
+42.85
++2.06
+Mixtral-8x22b-instruct 52.04
++1.53
+11.82
++1.82
+20.00
++0.00
+81.90
++2.61
+19.91
++1.42
+57.82
++2.06
+  • 1
+    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125
+Our evaluation  [^6] See the setting in Appendix N includes both closed-source LLMs and open-source LLMs.
 The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^7] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^8] https://deepmind.google/technologies/gemini/.
-The open-source LLMs include DeepSeek-v2 DeepSeek-AI (2024) and DeepSeek-coder Guo et al.
-(2024).
+The open-source LLMs include DeepSeek series by  DeepSeek-AI (2024),  Guo et al.
+(2024) and  Shao et al.
+(2024); Llama-3 models  [^9] https://github.com/meta-llama/llama3; Qwen-1.5 (Bai et al.) and Mixtral (Jiang et al.).
 Table 3 presents the evaluation results of different language models on the Mamo Benchmark.
-The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (indicated by the dagger symbol ).
-Table 3 displays the performance results when GPT-4-0613 is selected as the code modifier.
-Models ODE Optimization
-First order Second order System All Easy LP Complex LP
-GPT-4o †† 69.90% 74.55% 42.50% 68.21% 87.42% 29.86%
-GPT-4-turbo †† 67.85% 70.00% 42.50% 65.99% 88.19% 28.44%
-GPT-4 †† 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
-GPT-3.5-turbo †† 47.96% 60.91% 20.00% 48.84% 85.43% 10.43%
-Claude-3-Sonnet †† 58.16% 66.36% 35.00% 59.09% 85.12% 28.44%
-Claude-3-Haiku †† 53.06% 65.45% 25.00% 53.76% 86.81% 21.33%
-Gemini-1-pro †† 43.37% 54.55% 15.00% 43.64% 80.37% 16.27%
-Gemini-1.5-pro †† 67.35 % 67.27% 52.50% 65.61% 85.28% 34.12%
-DeepSeek-v2 †† 56.12% 68.18% 32.50% 57.23% 86.81% 6.16%
-DeepSeek-coder †† 33.67% 49.09% 17.50% 36.71% 78.83% 18.96%
-[…]
+The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (the improvement is indicated by the subscript).
+Table 5 in Appendix A shows results when GPT-4-0613 is selected as the code modifier.
 The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks.
-Specifically, GPT-4o achieves the highest performance in ODE tasks with 69.90% in first-order and 74.55% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 32.23% in complex LP scenarios when using the LLM to rectify syntax errors.
-Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with 48.27% overall in ODE tasks and 86.66% in easy LP tasks when enhanced with syntax correction, highlighting its potential despite being open-source.
+Specifically, GPT-4o achieves the highest performance in ODE tasks with 68.37% in first-order and 37.27% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 31.28% in complex LP scenarios when using itself as code modifier.
+Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability, achieving a 57.57% success rate across all tasks when enhanced with syntax correction by itself, underscoring its potential despite its open-source nature.
+Moreover, Mixtral-8x22B also exhibits competitive performance relative to DeepSeek-v2.
+After fix syntax error by itself, it achieves a remarkable score of 21.33% on the complex LP task, showcasing its potential in modeling complex optimization problems.
 The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models.
-For instance, GPT-3.5 Turbo’s performance in ODE tasks increased from 15.90% to 26.30% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 45.66% to 52.31%.
-Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 41.33% to 48.27% when enhanced with syntax correction.
-Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement.
-For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 48.27% with self-improvement to 57.23% when GPT-4 is used as the code modifier.
-This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
+For instance, GPT-3.5-turbo’s performance in ODE tasks increased from 12.72% to 18.21% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 32.95% to 38.73%.
+Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 32.08% to 35.55% when enhanced with syntax correction.
 6.3 On the Format Errors and Code Modifiers
-Models ODE Optimization
-format raw self-improved GPT-4-improved raw self-improved GPT-4-improved
-GPT-4o 96.82% 99.42% 99.42% 92.93% 97.91% 97.91%
-GPT-4-turbo 96.82% 99.13% 98.84% 93.86% 97.80% 97.56%
-GPT-4 94.51% 97.69% 97.69% 90.96% 95.13% 95.13%
-GPT-3.5-turbo 32.37% 60.40% 94.22% 87.95% 89.80% 96.76%
-Claude-3-Sonnet 84.68% 93.06% 97.40% 91.89% 93.97% 96.76%
-Claude-3-Haiku 78.03% 92.77% 98.27% 82.61% 92.58% 96.29%
-Gemini-1.0-pro 56.36% 69.94% 88.73% 88.06% 89.11% 94.21%
-Gemini-1.5-pro 93.64% 98.27% 98.84% 95.48% - 96.87%
-deepseek-v2 70.81% 83.81% 96.24% 90.27% 92.00% 94.90%
-deepseek-coder 88.44% 93.64% 95.95% 92.58% 96.29% 98.84%
-Table 4: Correct format rate on raw, self-improved, and GPT4-improved format.
-The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different improvement methods.
-The table considers three scenarios: no improvement, self-improvement, and improvement provided by GPT-4.
-The GPT-3.5-turbo model shows the most dramatic improvement when enhanced by GPT-4, jumping from a 32.37% success rate to 94.22%.
-Interestingly, Claude-3 models (Sonnet, and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4, highlighting the compatibility of GPT-4’s improvement methods with other models.
+Table 4: Execution Success Rates for Raw, Self-Modified, and GPT-4 Modified Methods, denoting scenarios without code modifiers, utilizing the LLM as a code modifier, and employing GPT-4-0613 as the code modifier.
+This table highlights the success rates of file execution (correct format) across these three scenarios, differing from the Table 3 which shows the correctness rate.
+Models ODE (%) Optimization (%)
+raw self-modified GPT-4-modified raw self-modified GPT-4-modified
+GPT-4o 96.82 99.42 99.42 92.93 97.91 97.91
+GPT-4-turbo 96.82 99.13 98.84 93.86 97.80 97.56
+GPT-4 94.51 97.69 97.69 90.96 95.13 95.13
+GPT-3.5-turbo 37.28 65.03 93.93 87.95 89.80 96.76
+Claude-3-Sonnet 84.68 93.06 97.40 91.89 93.97 96.76
+Claude-3-Haiku 78.03 92.77 98.27 82.61 92.58 96.29
+Claude-3-Opus 88.73 96.82 97.98 90.38 93.40 95.94
+Gemini-1.0-pro 56.36 69.94 88.73 88.06 89.11 94.21
+Gemini-1.5-pro 93.64 98.27 98.84 95.48 98.73 96.87
+DeepSeek-v2 70.81 83.82 96.24 90.27 92.00 94.90
+DeepSeek-coder 88.44 93.64 95.95 92.58 96.29 98.84
+DeepSeek-math-7B-base 28.90 29.77 79.77 55.50 58.40 75.67
+DeepSeek-math-7B-rl 5.20 6.64 80.64 19.93 21.55 55.97
+Llama-3-8B 20.23 26.01 68.21 75.09 76.36 79.14
+Llama-3-70B 24.56 30.92 77.46 77.64 79.49 89.46
+Qwen-1.5-72B-Chat 55.49 89.60 95.38 77.40 92.24 95.37
+Qwen-1.5-110B-Chat 51.45 63.29 95.09 71.26 84.01 90.15
+Mixtral-8x7B-instruct 50.87 69.08 91.62 84.13 87.83 96.29
+Mixtral-8x22B-instruct 80.92 88.73 96.24 87.72 93.51 96.29
+The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different modification methods.
+The table considers three scenarios: no modification, self-modification, and modification provided by GPT-4-0613.
+The GPT-3.5-turbo model shows the most dramatic improvement when modified by GPT-4, jumping from a 37.28% success rate to 93.93%.
+Interestingly, Claude-3 models (Sonnet and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4-0613, highlighting the compatibility of GPT-4’s modification methods with other models.
 In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
-The improvement from GPT-4 is generally superior to both self-improvement and no improvement.
+The modification from GPT-4 is generally superior to both self-modification and no modification.
 This indicates that GPT-4 can serve as a more effective and fair code modifier.
 By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors.
-Notably, in ODE tasks, the correct format rate for all models, except one, is generally greater than 94%, with the lowest being 88.73%.
-In Optimization tasks, the correct format rate is even higher, with the lowest rate being 94.21%.
+Table 5 in Appendix A shows the results of selecting GPT-4-0613 as code modifier.
+6.4 Analysis
+Given the results in Section 6.26.3 and Appendix A, several observations emerge from the benchmark results.
+For open-source models, larger models generally perform better, as evidenced by comparisons between Llama-3-70B and Llama-3-8B, Mixtral-8x22B-instruct and Mixtral-8x7B-instruct, with the larger models consistently achieving higher scores.
+From Table 4, it is apparent that some models struggle with adhering to the syntax of the solver, such as GPT-3.5-turbo and certain open-source models.
+Some models like Llama-3-8B may not closely follow the instructions, even worse than the syntax error.These models tend to show significant improvements in both correct rates (as indicated by the subscript in Table 3) and execution success rates after the application of code modifiers, particularly when their initial performance without modifiers is low.
+Notably, in the ODE part, GPT-3.5-turbo achieves a remarkable 28.03% increase in execution success rate by self-modified, accompanied by substantial improvements in ODE performance.
+In Optimization part, Qwen-1.5-72B-Chat follows the same trends.
+When comparing the results from Table 3 and Table 5, it becomes clear that the improvements are even more pronounced for open-source models when utilizing GPT-4 modifications rather than self-modifications.
+See Appendix E for the analysis of typical modeling errors and case study,
 7 Conclusion and Future Directions
-In this article, we introduce Mamo, a novel benchmark that utilizes solvers to evaluate the mathematical modeling capabilities of large language models (LLMs).
-We propose a shift in the evaluation paradigm from traditional outcome-oriented evaluation to process-oriented evaluation.
-This shift allows us to scrutinize the modeling process undertaken by LLMs, providing a more detailed and comprehensive assessment of their mathematical problem-solving abilities.
-Our evaluation focuses more on the mathematical modeling capabilities of the models, leaving the process of solving abstract mathematical problems to the solvers, thereby avoiding the impact of errors such as computational capabilities of the models on their actual modeling abilities.
-Our benchmark represents an important step in understanding the fundamental modeling capabilities of LLMs.
-By focusing on the modeling process rather than the correctness of the final solution, we are able to gain valuable insights into the strengths and limitations of current LLMs.
-This approach provides new directions for future research, emphasizing the importance of the modeling process in solving mathematical problems.
-The introduction of Mamo paves new avenues for future research.
-The benchmark can be expanded to include a wider variety of mathematical problems and solvers, broadening its applicability and providing a more comprehensive evaluation of the mathematical modeling capabilities of LLMs.
-Insights gained from the benchmark can be used to guide the development of future LLMs.
-By understanding the strengths and weaknesses of current models in solving mathematical problems, researchers can design new models that better handle these tasks.
-We will extend our model list to other models (Llama2 series, Mistral series etc) in our further study.
-Limitations
-While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research.
-The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning.
-This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities.
-Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
-Ethics Statement:
-The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes.
-Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups.
-This includes but is not limited to issues of privacy, security, and the potential for misuse of the model.
-Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.
-Appendix A Cross Review
+In conclusion, Mamo’s innovative benchmark shifts the evaluation focus of LLMs towards the modeling process, offering a nuanced assessment of their mathematical problem-solving capabilities.
+By isolating modeling from solving through the use of solvers, Mamo provides a comprehensive understanding of LLMs’ strengths and limitations in mathematical modeling.
+This approach not only enhances our insights into the fundamental modeling abilities of LLMs but also guides future research by highlighting the importance of the modeling process in solving mathematical problems.
+Appendix A GPT-4-0613 as Code Modifier
+[…]
+Models ODE Optimization Overall (%)
+First Second System Easy Complex
+order (%) order (%) (%) LP (%) LP (%)
+GPT-4o †† 68.37 37.27 40.00 87.42 29.38 68.07
+GPT-4-turbo †† 65.82 33.64 40.00 88.19 27.96 67.58
+GPT-4 †† 60.20 29.09 42.50 86.96 24.17 64.93
+GPT-3.5-turbo †† 44.90 20.00 20.00 85.43 9.48 57.48
+Claude-3-Opus †† 59.18 36.36 47.50 84.20 14.22 62.37
+Claude-3-Sonnet †† 56.63 26.36 32.50 85.12 27.96 63.44
+Claude-3-Haiku †† 51.53 25.45 22.50 86.81 18.96 61.54
+Gemini-1-pro †† 40.31 18.18 10.00 80.21 15.64 54.51
+Gemini-1.5-pro †† 65.82 30.91 50.00 85.28 34.12 67.08
+DeepSeek-v2 †† 54.08 30.91 30.00 86.81 3.79 60.05
+DeepSeek-coder †† 30.10 11.82 15.00 78.83 18.96 52.27
+DeepSeek-math-7b-base †† 13.78 9.09 5.00 59.05 9.48 36.72
+DeepSeek-math-7b-rl †† 13.78 7.27 7.50 47.70 1.90 29.20
+Llama-3-8b †† 11.22 0.91 0.00 63.19 8.53 37.47
+Llama-3-70b †† 14.80 7.27 10.00 70.71 9.48 43.18
+Qwen-1.5-72b-Chat †† 16.84 13.01 5.45 78.53 6.64 47.80
+Qwen-1.5-110b-Chat †† 40.82 13.64 20.00 78.83 14.22 53.51
+Mixtral-8x7b-instruct †† 32.14 11.82 20.00 75.46 9.95 49.38
+Mixtral-8x22b-instruct †† 55.10 17.27 22.50 85.43 22.27 61.21
+Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement.
+For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 32.08% with self-modification to 43.93% when GPT-4 is used as the code modifier.
+This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
+Appendix B Cross Review
 In ODE part, we have the following review process:
@@
 Finally, we have 346 problems in ODE part.
-We have also conducted the cross-review process. See the details in Appendix.  A.1 and the results on Appendix.  A.1.1.
+We have also conducted the cross-review process. See the details in Appendix B.1 and the results on Appendix B.1.1.
 In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions.
@@
 Among them, 8 of them were corrected and others (36 questions) were filtered out.
-In complex LP part, for the original 211 questions, 40 were found synthetized by wrong mathematical models.
+In complex LP part, for the original 211 questions, 40 were found formulated by wrong mathematical models.
 All of them were corrected.
-A.1 Cross-Review Process
+B.1 Cross-Review Process
 Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis.
-Four independent reviewers (see their qualification in Section B) were tasked with assessing 50 questions each.
+Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^10] The hourly wage pay is 100 RMB/h, the total amount spent on the participant compensation is 2400 RMB, approximately 331 dollars.
 The responses could either be numeric or labeled as “error” if a question was deemed problematic.
@@
 Here are our results:
-A.1.1 Results
+B.1.1 Results
 The effectiveness of the review process was quantified using the metrics defined in Subsection 6.1.2, focusing on the accuracy and inter-rater reliability among reviewers.
@@
 These results provide insights into the overall effectiveness and areas of improvement for the cross-review process, particularly in terms of aligning understanding and interpretation of the evaluation criteria among reviewers.
-Appendix B Annotator Qualifications
+Appendix C Annotator Qualifications
 The benchmarking process benefits from the expertise of reviewers with robust backgrounds in mathematics, underscored by their academic accomplishments.
@@
 degrees in mathematics, further solidifying the depth of expertise and analytical skills available for the benchmarking process.
-B.1 Qualification of Reviewers
-We select the reviewers mentioned in Appendix  A.1 based on the following criteria:
+C.1 Qualification of Reviewers
+We select the reviewers mentioned in Appendix B.1 based on the following criteria:
   1. 1.
@@
 These criteria ensure their ability to read questions in English and solve ODE problems.
-Appendix C Few-shot Experiment
+Appendix D Metric
+The comparison process is as the following:
+$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (1)} or }|O-A|<1\times min\{10^{-n},10^{-2}\}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (2)},}\\ 0&\text{otherwise.}\end{cases}$
+Our consideration: (1) The difference is much less than the Answers, so can be omitted.
+(2) The only difference lies in the last few digits of the required significant figure of the answer.
+By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|O-A|<\xi\times min\{10^{-n},10^{-2}\}$.
+We have explored the relationship between various thresholds and the accuracy of GPT-4o.
+Table 6: Accuracy at various thresholds for different categories(%), where $\xi_{i}=i$
+$\xi$ 0 1 2 3 4 5 6 7 8 9
+ODE 48.55 50.87 54.62 56.36 57.51 58.96 59.83 60.12 60.69 61.85
+Easy LP 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27
+Complex LP 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75
+Figure 2: Accuracy of different thresholds in different categories
+Figure 2 plots the results. In our metric, we choose strict $\xi=1$.
+Appendix E Error Analysis
+In this section, we will analyze some typical errors.
+E.1 Deviation from Instruction
+It is common for some models that show a low capability of few-shot learning, especially for small models.
+For example, in a case Llama-3-8B output ‘I wish you good luck’ when prompting with some few-shot demonstrations in mathematical modeling.
+E.2 Syntax Error
+In ODE parts, the syntax errors is obviously the syntax errors in the reponsed python code.
+In Optimization part, the .lp file require more strict syntax than the universal optimization formulation.
+The Figure 4 and 4 show the typical syntax errors in .lp format.
+Figure 3: The syntax error is caused by the constraint, in .lp format, the variables should be placed on the left hand side of the inequality
+Figure 4: Same type of syntax error as Figure 4
+E.3 Modeling Error
+One of the most common errors in modeling is the lack of recognition of decision variable types.
+The types of decision variables in optimization problems are divided into discrete (sometimes integer, sometimes binary) and continuous.
+Figure 5 is an example.
+Figure 5: In this TSP problem, the decision variable should be binary, instead of "General", which means “integer”
+Appendix F Scope of Mamo
+Table 7: The taxonomy of mathematical modeling. In the Mamo benchmark, we only include the categories where sophisticated solvers are available.
+Category In Mamo
+Changes and Differences
+Proportionality and Geometric Similarity
+Optimization Problem Modeling
+Differential Equations
+Probabilistic Modeling
+The scope of Mamo is chosen under the consideration in Section 4.2. As Table 7 shows, Mamo focus on the ODE and Optimization areas.
+Appendix G Examples of Different ${\widehat{M}}$
+Different models will typically lead to different final answers after solving by the solvers. For example, in ODE part:
+${\widehat{M}}_{1}:y^{\prime}+y=0$
+and
+${\widehat{M}}_{2}:y^{\prime}+2y=0$
+Suppose we ask for $y(1)$, then we have ${\widehat{A}}_{1}=e^{-1}$ and ${\widehat{A}}_{1}=e^{-2}$,they have different values.
+Then for the optimization,
+${\widehat{M}}_{1}$:
+$\max_{x,y}$ $x+y$
+subject to $x+y\leq 20,$
+$x-y\geq 0,$
+$0\leq x\leq 10,$
+$0\leq y\leq 5,$
+$x,y\in\mathbb{Z}.$
+${\widehat{M}}_{2}$:
+$\max_{x,y}$ $x-y$
+subject to $x+y\leq 20,$
+$x-y\geq 0,$
+$0\leq x\leq 10,$
+$0\leq y\leq 5,$
+$x,y\in\mathbb{Z}.$
+the optimal value of ${\widehat{M}}_{1}$ is ${\widehat{A}}_{1}=10$, while the optimal value of ${\widehat{M}}_{2}$ is ${\widehat{A}}_{2}=0$.
+The minor difference in the mathematical models will result in different final values.
+Appendix H Examples of Syntax Errors
+Some outputs of LLM contain the correct model but not follow the syntax of solvers. For examples, the standard .lp file for
+$\max_{x,y}$ $x+y$
+subject to $x+y\leq 20,$
+$x-y\geq 0,$
+$0\leq x\leq 10,$
+$0\leq y\leq 5,$
+$x,y\in\mathbb{Z}.$
+is
+Minimize
+obj: x + y
+Subject To
+    c1: x + y <= 20
+    c2: x - y >= 0
+Bounds
+    0 <= x <= 10
+    0 <= y <= 5
+Generals
+    x
+    y
+End
+However, if we replace the constraint $c1$ with x <= 20 - y, the code will be reported as an error, even if the two constraints are equivalent.
+To specifically test the modelling ability of the LLMs, we apply code modifier to fix the syntax errors.
+Appendix I Few-shot Experiment
 To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance.
 In our analysis, we tested the performances of 11 models on ODE problems using prompts with 0 shot, 1 shot, 3 shots, 5 shots, and 10 shots.
-The results, shown in Figure  3, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots.
+The results, shown in Figure 6, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots.
 Similarly, we performed sensitivity tests on optimization problems, assessing the accuracy of 10 models with prompts of 0 shot, 1 shot, 3 shots, and 5 shots.
-The results are depicted in Figure  4.
+The results are depicted in Figure 7.
 Figure 6: Results of models accuracy and shots on ODE problems
@@
 The accuracy of GPT-4-0613 was 66.56% in Easy LP and 14.69% in Complex LP, resulting in an overall LP accuracy of 53.65%, consistent with the results obtained from sampling.
-Appendix D Testing Process
-We use 3-shot learnging in testing, see examples in  G
+Appendix J Testing Process
+We use 3-shot learning in testing; see examples in Appendix O. The examples of testing process in ODE and optimization is shown in Figure 8 and Figure 9.
 Figure 8: Example of testing process in optimization
 Figure 9: Example of testing process in ODE
-Appendix E Data Synthesis
-Figure  7 shows the example of synthesis process in ODE.
+Appendix K Question Formulation
+Figure  10 shows the example of synthesis process in ODE.
 Figure 10: Example of synthesis process in ODE
-Appendix F Evaluation script
+Appendix L Word Cloud
+See the word cloud of the combined data in Figure 11
+Figure 11: The word cloud of the data after filtering the requirement words such as ’significant figures’, ’rounded’
+Appendix M Evaluation script
 def comp(output, standard_answer):
@@
         rate = dif / float(standard_answer)
-    if abs(rate) <= 1e-4:
+    if abs(rate) < 1e-4:
         return 1
@@
         digit = len(standard_answer.split(’.’)[1])
+        if digit <= 2:
+            digit = 2
         s_ans = float(standard_answer) * 10 ** digit
         ans = float_output * 10 ** digit
-        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
+        return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
     else:
-        s_ans = float(standard_answer)
-        ans = float_output
-        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
-Appendix G Prompts
-Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems.
+        digit = 2
+        s_ans = float(standard_answer) * 10 ** digit
+        ans = float_output * 10 ** digit
+        return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
+Appendix N Settings
+We have the following settings:
+  • We testing the following models by calling API: All the closed source models, DeepSeek-V2 and DeepSeek-coder , with temperature=0.8
+  • Other models was deployed and inference using VLLM (Kwon et al., 2023) on machine in slurm cluster with 8 * A100, with temperature 0.8 and max_length 4096 tokens.
+Appendix O Prompts
+[Instruction] Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems.
 Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format.
@@
 - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables.
-Here comes the examples: Example_1: (the input) A manufacturing company produces two types of products: $X$ and $Y$.
+Here comes the examples: [Example_1] (the input) A manufacturing company produces two types of products: $X$ and $Y$.
 The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$.
@@
 Figure 12: The prompt for testing in optimization part. [question] refers to a question in the benchmark. […] refers to the examples. We use 3-shot in testing
-Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy.
+[Instruction] Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy.
 Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions.
@@
 Use Python’s ‘round‘ function for decimal rounding.
-For significant figures, include and use the following function in your code: ```$python$ def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 `````````` ` ` 4.
+For significant figures, include and use the following function in your code: “‘ python def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 “‘ 4.
 Your response should be entirely in Python code, formatted to run directly without modifications.
@@
 Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution.
-Here comes the examples: […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements.
+Here comes the [examples] : […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements.
 And Please directly generate the code without any explaintion(except the comments in the code).
-the following lines are forbidden: ′′′ Here’s the Python code that solves the given problem and meets the specified requirements: `````````` ` `python `````````` ` ` This code uses the `$solve_{i}vp`$ function from SciPy to solve the initial value problem for the given differential equation.
-The `$round_{t}o_{s}ignificant_{f}igures`$ function is included to round the final answer to the specified number of significant figures.
+the following lines are forbidden: Here’s the Python code that solves the given problem and meets the specified requirements: “‘ python “‘ This code uses the ‘solve_ivp’ function from SciPy to solve the initial value problem for the given differential equation.
+The ‘round_to_significant_figures’ function is included to round the final answer to the specified number of significant figures.
 Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures.
-′′′ Please avoid the above sentences.
+Please avoid the above sentences.
 Take a deep breathe before answering the question.
@@
 Figure 13: The prompt for testing in ODE part. [question] refers to a question in the benchmark. […] refers to the examples.
-You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully.
+[Instruction] You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully.
 If there is no error, please return the original code.
@@
 PLEASE do not response anything except the code, No other comments outside the code are allowed.
-The followings are forbidden: ”’ The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘.
+The followings are forbidden: The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘.
 The round function is missing a closing parenthesis.
-Here is the corrected version: “‘python “‘ ”’ Please avoid resposing above sentences.
+Here is the corrected version: “‘ python “‘ Please avoid resposing above sentences.
 Please output only the corrected code, with no additional text or explanations.
@@
 Figure 14: The prompt for fixing syntax error in ODE. [error] refers to the message when executing the Python code. While [code] refer to the code which contains syntax error
-You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully.
+[Instruction] You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully.
 If there is no error, please return the original code.
@@
 PLEASE do not response anything except the code, No comments are allowed.
-The followings are forbidden: ”’ This is the correct .lp file: “‘lp “‘ ”’ Please avoid resposing above sentences.
+The followings are forbidden: This is the correct .lp file: “‘ lp “‘ Please avoid resposing above sentences.
 Please output only the corrected code, with no additional text or explanations.
@@
 Figure 15: The prompt for fixing syntax error in LP. [lp_code] refer to the lp code which contains syntax error
+Appendix P Limitations
+While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research.
+The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning.
+This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities.
+Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
+Appendix Q Ethics Statement:
+The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes.
+Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups.
+This includes but is not limited to issues of privacy, security, and the potential for misuse of the model.
+Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.

### v3

@@
-Mamo: a Mathematical Modeling Benchmark with Solvers
+LLMs for Mathematical Modeling: Towards Bridging the Gap between Natural and Mathematical Languages
 Abstract
-Mathematical modeling involves representing real-world phenomena, systems, or problems using mathematical expressions and equations to analyze, understand, and predict their behavior.
-Given that this process typically requires experienced experts, there is an interest in exploring whether Large Language Models (LLMs) can undertake mathematical modeling to potentially decrease human labor.
-To evaluate of LLMs in mathematical modeling, we introduce a new benchmark, Mamo, that transcends traditional result-oriented assessments.
-Unlike conventional methods that primarily assess LLMs based on the accuracy of solutions to mathematical problems, our approach offers deeper insight into the modeling process itself.
-By focusing on the processes LLMs undertake rather than the correctness of their final solutions, Mamo pioneers a novel evaluation paradigm.
-This shift underscores the importance of understanding the inherent modeling capabilities of LLMs, paving the way for a more nuanced and comprehensive analysis of their problem-solving strategies.
-Our work marks a significant advancement in the field, suggesting a new direction for future research by emphasizing the evaluation of LLMs’ modeling processes over the mere correctness of answers.
-This benchmark not only facilitates a better understanding of LLMs’ mathematical modeling capabilities but also sets a new standard for evaluating their performance in complex problem-solving scenarios.
+Large Language Models (LLMs) have demonstrated strong performance across various natural language processing tasks, yet their proficiency in mathematical reasoning remains a key challenge.
+Addressing the gap between natural and mathematical language requires advanced reasoning capabilities, approaching those of Artificial General Intelligence (AGI).
+However, the evaluation remains challenging, as perfectly representing reality is inherently elusive, and traditional methods like manual or direct comparison of mathematical statements (Ramamonjison et al., 2023) are insufficient for assessing true modeling ability.
+We propose a process-oriented framework to evaluate LLMs’ ability to construct mathematical models, using solvers to compare outputs with ground truth.
+Introducing Mamo, a benchmark with 1,209 questions covering ordinary differential equations, linear programming, and mixed-integer linear programming, we enable automatic evaluation of modeling accuracy.
+The results show that existing LLMs struggle with complex mathematical modeling tasks, with larger models demonstrating superior performance, while open-source models remain competitive in simpler cases but still fall short of proprietary models in more challenging problems.
+LLMs for Mathematical Modeling: Towards Bridging the Gap between Natural and Mathematical Languages
+Xuhan Huang1, Qingning Shen1, Yan Hu1thanks: Corresponding author: huyan@cuhk.edu.cn, Anningzhe Gao2, Benyou Wang1 1 The Chinese University of Hong Kong, Shenzhen 2 Shenzhen Research Institute of Big Data xuhanhuang,qingningshen@link.cuhk.edu.cn, huyan@cuhk.edu.cn anningzhegao@gmail.com, wangbenyou@cuhk.edu.cn
 1 Introduction
-Recent advancements in Large Language Models have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al.
-(2022); OpenAI (2023); Nijkamp et al.
-(2022); Tang et al.
-(2024).
-However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al.
-(2022).
-This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
-Natural language captures human complexity and nuance, while mathematical language excels in precision and modeling the natural world (Giordano et al., 2013).
+Recent advancements in Large Language Models (LLMs) have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks (Ouyang et al., 2022; OpenAI, 2023; Nijkamp et al., 2022; Tang et al., 2024).
+However, the domain of mathematical reasoning remains a critical aspect of the LLM evaluation (Wei et al., 2022).
+This focus gauges LLMs’ proficiency in solving mathematical challenges and reveals their underlying abilities in abstract reasoning and logic.
+Natural language reflects the complexity and subtlety of human communication, while mathematical language provides accuracy and is ideal for representing the natural world (Giordano et al., 2013).
 Bridging their gap requires high-level cognitive skills akin to Artificial General Intelligence (AGI), a task at which LLMs show promise.
-However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
-In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem.
-They search through possible options to maximize or minimize an objective under given constraints if any.
-By using solvers, The challenge becomes problem formulation under a given solver or a solver set.
-becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers.
-Recent studies have already combined LLMs with solvers Yang et al.
-(2023); Feng et al.
-(2023); Pan et al.
-(2023).
+Recent studies have already combined LLMs with solvers (Feng et al., 2023; Pan et al., 2023).
 In the field of optimization, AhmadiTeshnizi et al.
-(2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems.
+(2023) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems.
 The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver.
 This implies the potential for testing LLM’s modeling capabilities.
-We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of LLMs.
-Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies.
-We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
-The contributions of this paper are as below: (1) We introduce a new benchmarking strategy to assess mathematical modeling through Solvers, facilitating a rigorous evaluation process.
-(2) We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities.
-Mamo encompasses a wide array of modeling questions (With a total of 1209 meticulously curated questions), including ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks.
-2 Related Work
-Recent studies have made significant strides in applying LLMs to mathematical problem-solving, introducing innovative datasets for comprehensive evaluation.
-Frieder et al.
-(2024) and Yuan et al.
-(2023) have focused on datasets that test LLMs across various math problems and arithmetic expressions, respectively.
-Lewkowycz et al.
-(2022) explored training models using natural language paired with LaTeX-formatted mathematical content from arXiv, enhancing syntactic understanding.
-Zhang et al.
-(2024) developed the CARP dataset for computation-intensive challenges, while He et al.
-(2024) introduced OlympiadBench, a bilingual and multimodal benchmark for competition-level mathematical reasoning and proofs.
-Liu et al.
-(2024) constructed MathBench, which is designed with a carefully structured difficulty hierarchy and focuses on assessing the understanding of theoretical knowledge in LLMs.
-AhmadiTeshnizi et al.
-(2023a) constructed NLP4LP, a benchmark comprising 52 linear programming (LP) and mixed-integer linear programming (MILP) problems.
-These efforts collectively push the boundaries of LLMs in mathematical cognition, showcasing the growing complexity and specificity of tasks LLMs are being trained to tackle.
-The pursuit of enhancing and devising novel solvers has seen a variety of research efforts, with a notable focus on the integration of LLMs.
-Yang et al.
-(2023) investigated the potential of employing LLMs as direct solvers, pioneering the exploration of LLMs’ autonomous problem-solving capabilities.
-He-Yueya et al.
-(2023) introduced a hybrid approach that marries an LLM with an external symbolic solver, specifically for equation solving, marking a significant advancement in combining neural and symbolic computation.
-Furthermore, Pan et al.
-(2023) developed LOGIC-LM, a cutting-edge framework that fuses LLMs with symbolic solvers, aimed at enhancing the solving of logical problems.
-In a closely related study, AhmadiTeshnizi et al.
-(2023a) introduced OptiMUS, an LLM-based agent explicitly crafted for formulating and solving optimization problems, thus demonstrating the broad applicability and potential of LLMs in tackling intricate computational challenges.
-These endeavors underscore the innovative directions being explored in the realm of solver development, leveraging the robust capabilities of LLMs.
-3 Background
-3.1 Mathematical Modeling
-Giordano et al.
-(2013) emphasize the role of mathematical models as intermediaries that transform real-world problems into structured mathematical forms, facilitating the discovery of solutions.
-Despite LLMs’ proven competencies across various tasks, their capacity for developing mathematical models remains uncertain.
-This involves two key phases: model formulation, which demands a thorough comprehension of the problem, and model resolution, typically addressed by computational solvers.
-To accurately gauge LLMs’ potential in this domain, our strategy focuses on their ability to construct mathematical models, entrusting calculations to specialized solvers.
-This approach highlights LLMs’ creativity and generalization skills, simultaneously capitalizing on solvers’ computational efficiency.
-Motivationsto Benchmark Mathematical Modeling for LLMs   We evaluate mathematical models for two reasons.
-The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can.
-Mathematical modeling is considered a highly advanced capability that we believe experts possess.
-This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence.
-On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al.
-(2021), etc., show that the performance has become saturated.
-Even a 7B model Yu et al.
-(2023b); Liu et al.
-(2023) (sometimes with a verifier-like model Yu et al.
-(2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4.
-Therefore, we believe there is a need to construct more challenging mathematical datasets.
-Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
-3.2 Using Solvers in Mathematical Modeling
-The model generated by the LLM is then subjected to verification against real-world data to assess its validity, ensuring it addresses the problem logically and sensibly.
-The solver, an algorithmic tool designed to find solutions to mathematical models, is also involved in this process and it is also beneficial for verification.
-Different types of solvers are specialized for various classes of problems, such as optimization solvers.
-However, they share the common property: They are all algorithmic tools designed to find solutions to mathematical models.
-Rationale of Using solver in Mathematical Modeling for LLMs These solvers operationalize the abstract constructs created by LLMs, translating them into tangible outputs that can be analyzed and validated against real-world data and phenomena.
-The solver is a crucial instrument that resolves the model presented by the LLM, and by comparing the solution it provides with the correct answer, we assess the accuracy of the mathematical model generated by the LLM.
-4 Philosophy of Mamo Benchmark
-4.1 Methodology: exact answer verification with solvers
-We utilize solver to resolve this mathematical model.
-The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps.
-By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$), since typically the wrong ${\widehat{M}}$ will result in wrong ($\widehat{A}$), like the examples in Appendix  G.
-This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
+However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
+To address this, we propose a new benchmarking strategy that evaluates mathematical modeling by utilizing solvers and answer verification.
+The core philosophy is to use solvers to resolve mathematical models and assess their accuracy by comparing the solver’s output with the ground truth.
+Additionally, we introduce Mamo, a novel benchmark designed to assess the mathematical modeling capabilities of LLMs, leveraging solvers to shift the focus from conventional outcome-based evaluations to a more process-oriented approach.
+This strategy is particularly rational because it allows us to isolate and scrutinize the entire modeling process, from formulation to solution, providing a more holistic and rigorous evaluation of an LLM’s mathematical reasoning abilities.
+The contributions of this paper are as follows:
+  • We introduce an automatic evaluation framework to assess mathematical modeling through solvers, enabling exact answer matching.
+  • We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities.
+Mamo encompasses a wide array of modeling questions (with a total of 1,209 meticulously curated questions), including ordinary differential equations (ODEs) and optimization problems within linear programming and mixed-integer linear programming frameworks.
+2 An Evaluation Framework for Mathematical Modeling
+2.1 Definition: Mathematical Modeling
+Mathematical modeling is widely considered an advanced skill typically possessed by experts.
+According to Giordano et al.
+(2013), mathematical models act as intermediaries that convert real-world problems into structured mathematical forms, facilitating the discovery of solutions.
+As illustrated in Figure 1, we aim to map a natural language question $Q$ into a mathematical model $M$, which is expressed in mathematical language as follows:
+$f:Q\rightarrow M.$
+By solving this model, we can obtain an answer $A$ to the original problem $Q$.
+Traditional mathematical modeling is a labor-intensive, highly specialized process that depends heavily on human expertise.
+2.2 Gap between Natural and Mathematical Languages
+Natural language captures the complexity and nuance of human thought, while mathematical language excels in precision and in modeling natural phenomena (Giordano et al., 2013).
+A single statement in natural language may correspond to different mathematical representations, and conversely, distinct natural language statements may map to the same mathematical formulation.
+For example, consider the natural language statement: The car starts with an initial velocity of zero. This can correspond to at least two different mathematical representations:
+Mathematical formulation 1: $y^{\prime}(0)=0$
+Mathematical formulation 2: $v(0)=0$.
+Conversely, these two mathematical expressions may also correspond to natural language 2: The population growth rate is initially zero.
+Bridging the gap between natural language and mathematical language requires advanced cognitive abilities, a task where LLMs have shown significant potential.
+However, evaluating the effectiveness of mathematical modeling remains challenging, as achieving a perfect representation of reality through models is often unattainable.
+Figure 1: $Q$ is the natural language problem, while $M_{1}$,$M_{2}$ are different mathematical models.
+Challenges in Evaluation
+Due to the complex mapping between natural language and mathematical language, a real-world problem $Q$ may correspond to multiple equivalent mathematical models, such as $M_{1},M_{2},M_{3},\dots$.
+As illustrated in Figure 1, for a given problem $Q$, we have two distinct mathematical models, $M_{1}$ and $M_{2}$, both of which correctly describe the same real-world problem and are considered equivalent in the context of $Q$.
+Traditionally, the evaluation of mathematical models, including determining their equivalence, has been performed by human experts, a process that is both labor-intensive and time-consuming.
+In optimization modeling area, Ramamonjison et al.
+(2023) directly compare mathematical statements by standardizing variable names and evaluating accuracy based on declaration matches.
+While this approach addresses differences in variable naming, the abstraction of variables from natural language context, which is critical in mathematical modeling, was not examined.
 Figure 2: The pipeline to use exact answer verification via an additional solver.
-4.2 Scope of Mathematical Modeling
-We construct the category of Mathematical Modeling by referring to Giordano et al.
-(2013).
-In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Table 7 in Appendix F.
-Given the broad spectrum of mathematical modeling and the availability of specific solvers, our benchmark is meticulously designed to encompass domains where these tools can be effectively utilized.
-This section outlines the rationale for our focus on certain areas of mathematical modeling, based on solver availability and the clarity of the modeling process.
-Availability of Sophisticated Solvers: Advanced optimization solvers, such as COPT Ge et al.
-(2023), along with Python libraries for solving ordinary differential equations (ODEs), offer a strong foundation for testing the capabilities of Large Language Models (LLMs).
+2.3 Evaluating Modeling Using Solvers and Answer Matching
+The Evaluation Framework
+To facilitate the automatic evaluation of mathematical models, we propose an evaluation framework for mathematical modeling based on exact answer verification using solvers, as illustrated in Figure 2.
+A solver is an algorithmic tools designed to find solutions to various classes of problems.
+For instance, Gurobi is a well-known optimization solver (Gurobi Optimization, LLC, 2024), while Python packages (Virtanen et al., 2020) and MATLAB (Inc., 2022) are commonly used for solving ODEs.
+Solvers are essential for operationalizing abstract models, denoted as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$, generated by large language models (LLMs), translating them into concrete solutions $\widehat{A}$ that can be analyzed and validated against real-world data.
+For a given real-world problem and its corresponding answer pair $(Q,A)$, LLMs generate a mathematical model, denoted ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$.
+The solver resolves this model, producing a solution $\widehat{A}$, which is then compared to the correct answer $A$.
+This framework directly evaluates the LLM’s problem-solving capability, with any inaccuracies in $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$ typically leading to incorrect solutions $\widehat{A}$.
+By focusing on the final output, we avoid additional computational steps that could obscure the evaluation of the LLM’s performance.
+The Philosophy
+The philosophy behind our evaluation framework is goal-driven, based on the idea that the ultimate aim of mathematical modeling is to produce accurate solutions to real-world problems.
+We prioritize the correctness of the solution $\widehat{A}$ over the intermediate steps of model formulation.
+By comparing the solver’s output with the correct answer, we evaluate the effectiveness of the LLM’s generated model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$.
+The assumption here is that if the model is well-constructed, it will lead to an accurate solution.
+Despite its goal-driven structure, the framework retains a process-oriented emphasis on the modeling capability of LLMs.
+Unlike traditional benchmarks that treat answer correctness as an isolated metric, our approach explicitly evaluates the intermediate step of model formulation.
+The evaluative focus lies on whether the LLM constructs a mathematical model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$ that faithfully translates the real-world problem into mathematical language.
+Thus, while the solver’s answer validates the outcome, the core objective remains to assess the modeling process.
+This dual perspective allows us to maintain a scalable yet principled evaluation framework.
+While the solver’s output remains the validation mechanism, the primary objective is to assess the LLM’s ability to conceptualize and formulate mathematical problems rather than simply solving predefined mathematical tasks (more discussions are provided in Appendix M).
+3 Mamo Benchmark
+To implement this framework, we developed a new benchmark, Mamo, specifically for mathematical modeling.
+3.1 Scope of Mamo Benchmark
+Given the broad spectrum of mathematical modeling Giordano et al.
+(2013) (see Appendix G) and the availability of specific solvers, our benchmark is meticulously designed to encompass domains of ODEs and linear programming problems.
+This section outlines the rationale behind our focus on these areas, driven by the criteria of solver availability and result verifiability.
+Availability of Sophisticated Solvers:
+Advanced optimization solvers, such as COPT (Ge et al., 2023), along with Python libraries for solving ODEs, offer a strong foundation for testing the capabilities of LLMs.
 These tools enable a detailed and automated evaluation of the LLMs’ abilities to abstract, formulate, and accurately solve mathematical problems within these specific domains.
-Challenges in Modeling Changes and Differences, Proportionality and Geometric Similarity: Although simpler tools like calculators could address problems related to modeling changes and differences, proportionality and geometric similarity, the lack of a defined, structured modeling process complicates the implementation of our final-state-approach benchmark.
-Our objective is to assess the LLMs’ skill in developing comprehensive mathematical models that adhere to a solvable framework.
-The absence of a distinct modeling procedure in these areas impedes a standardized and objective assessment of an LLM’s modeling efficiency.
-Following the considerations mentioned above, we have developed our benchmark in the fields of optimization, and ODE, due to the lack of advanced and universal Partial Differential Equations (PDE) solvers.
-In the optimization segment, our focus is on Linear Programming (LP) and Mixed-Integer Linear Programming (MILP) problems.
+The Straightforward Verifiability of Results
+While calculators address basic problems in these areas, the absence of a structured modeling approach hinders the assessment of LLMs’ ability to create solvable mathematical models.
+Building on these considerations, we developed a benchmark focused on optimization and ODE, given the lack of universal PDE solvers.
 Compared to NLP4LP, introduced by AhmadiTeshnizi et al.
-(2023b) and focusing exclusively on optimization, our benchmark includes a larger number of problems.
-This benchmark set is highly scalable, with potential future expansions into nonlinear optimization and probabilistic modeling.
-Additionally, as new and advanced solvers emerge, it could be expanded into areas like Partial Differential Equations (PDE) and beyond.
-5 Data Collection
-5.1 Data Selection and Source Credibility
-The dataset for our benchmark consists of a blend of manually selected and GPT-synthetized questions.
-All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance.
-The data collection and synthesis process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines  [^2] See the details of the annotator qualifications in Appendix C.
-Manually Selected Questions
-The manually selected questions are high quality and reliable.
-(1)Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010).
-The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario.
-The source ensure the correctness of the answer of each question.
-(2) Transformation: Each question is meticulously transformed to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology.
+(2023) and focusing exclusively on optimization, our benchmark includes a larger number of problems, with potential expansions into nonlinear optimization, probabilistic modeling, and PDE as new solvers emerge.
+3.2 Data Selection
+Our benchmark dataset comprises manually selected and GPT-generated questions, curated by a team with advanced mathematical education, including Ph.D.
+holders, and an average GPA of 3.89/4.0, ensuring strong analytical expertise.
+Details on the collectors’ qualifications are provided in Appendix C.
+Source Reliability
+For manual selection, we referenced textbooks and solutions in ODE and optimization (Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce and DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Bertsimas and Tsitsiklis, 1997; Hurlbert, 2010).
+We focused on deriving mathematical models and assigning realistic scenarios, ensuring answer accuracy.
+Each question was adapted to meet the benchmark criteria (Section 3.3) by adjusting scenarios or solutions as needed.
 Question Formulation and Verification
-The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity.
-This is followed by the computation of answers to establish a ground truth.
-Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems.
-This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations.
-See the example in Appendix K.
-5.2 Data Quality Checking
-In the ODE part, we employed GPT-4 to confirm the solvability of newly generated questions, filtering out invalid ones.
-A thorough evaluation identified questions lacking information, with incorrect answers, and unclear statements.
-Consequently, problematic questions were deleted or corrected, resulting in a refined dataset.
-In the optimization section, particularly with the easy LP part, invalid questions were filtered out, and those with incorrect answers or misleading descriptions were corrected.
-In the complex LP part, questions formulated with wrong mathematical models were corrected.
-The summary of the review process is presented in Table 1.
-Category Initial Questions Filtered Final Questions
-ODE Part 383 37 346
-Easy LP Part 688 48 640
-Complex LP Part 211 0 211
+Data synthesis starts by creating mathematical constructs with random parameters, followed by answer computation to establish ground truth.
+GPT-4 then generates real-life scenarios, translating these models into context-rich problems, testing both mathematical reasoning and real-world contextualization.
+See the example in Appendix N.
+3.3 Guidance of Data Collection
+Figure 3: $Q$ is the natural language problem (may be ODE problem or optimization problem), where y is a solution function in ODE problems.
+Category Initial Questions Filtered Corrected Final Questions
+ODE 383 37 41 346
+Easy LP 688 36   8 652
+Complex LP 211   0 40 211
 Table 1: Summary of the review process.
-Reviewer Accuracy Rate
-Reviewer 1 90.0%
-Reviewer 2 84.0%
-Reviewer 3 88.0%
-Reviewer 4 74.0%
-Table 2: Accuracy of reviewers.
-For the cross-review process, 50 questions from the ODE dataset were selected for deeper analysis by four independent reviewers.
-Reviewers evaluated the questions, marking responses as either numeric answers or "error" if the question was problematic.
-The effectiveness of this review was quantified using metrics like Cohen’s Kappa to measure inter-rater reliability and accuracy rates for individual reviewers.
-The average Cohen’s Kappa was 0.60, indicating moderate to substantial agreement among reviewers, with individual Kappa values ranging from 0.50 to 0.71.
-Accuracy rates are moderately high as shown in Tables 2.
-These results highlight both the effectiveness of the review process and areas where alignment and training can improve consistency and accuracy among reviewers.
-See the review details in Appendix B.
-5.3 Guidance of Data Collection
-The collection of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs).
-The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification.
-The following are the key considerations in our benchmark guidance:
+Our benchmark dataset follows specific criteria to ensure the problems are suitable for automated verification.
+Figure 3 provides examples of these criteria, with further explanations presented in Appendix T.
 Property 1.
-Final-State Approach Solvability  Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
+Final-State Approach Solvability  Problems should be structured so that a solver can directly compute the solution once the LLM formulates the mathematical model.
+For example, asking for the optimal value of an optimization problem is solvable using a ’final-state approach’, whereas transforming an ODE solution requires extra steps and is not solvable this way.
 Property 2.
-Unified and Numerical Answers: To handle potentially various valid solutions for one problem: For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$.
-For optimization problems, questions should aim to find the optimal value.
-These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem.
-Besides, the answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
+Unified and Numerical Answers
+For ODEs, questions should seek the function’s value at a specific time $t_{0}$, and for optimization problems, the optimal value, ensuring a unified and numerical answer for clear comparison with the solver’s output.
 Property 3.
-Significant Figures and Precision  Each question will explicitly state the required significant figures or the level of numerical precision for the answer.
-This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
+Significant Figures and Precision
+Each question will explicitly state the required significant figures or the level of numerical precision for the answer.
 To assess the LLM’s capability to abstract mathematical models from real-world scenarios, we introduce Property 4.
 Property 4.
-Real-World Problem Context  Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
+Real-World Problem Context Questions should be framed as real-world problems without explicitly stating the underlying mathematical model.
+Reviewer Reviewer 1 Reviewer 2 Reviewer 3 Reviewer 4
+Reviewer 1 - 0.67 0.71 0.56
+Reviewer 2 0.67 - 0.65 0.50
+Reviewer 3 0.71 0.65 - 0.52
+Reviewer 4 0.56 0.50 0.52 -
+Table 2: Inter-reviewer Cohen’s Kappa scores.
+Reviewer Accuracy Rate (%)
+Reviewer 1 90.0
+Reviewer 2 84.0
+Reviewer 3 88.0
+Reviewer 4 74.0
+Table 3: Accuracy of reviewers.
+3.4 Data Quality Checking
+In the ODE section, we utilized GPT-4 to verify the solvability of generated questions as stated in Property 1, filtering out any invalid questions.
+A detailed evaluation identified issues such as missing information, incorrect answers, and unclear phrasing, leading to the removal or correction of problematic questions and resulting in a refined dataset.
+In the optimization section, specifically in the Easy_LP part, invalid questions were eliminated, and those with incorrect answers or misleading descriptions were corrected.
+For the Complex_LP section, questions involving incorrect mathematical models were revised (see an example in Appendix J).
+A summary of the review process is provided in Table 1.
+For the cross-review process, 50 questions from the ODE dataset were selected for in-depth analysis by four independent reviewers  [^1] The hourly wage is approximately 20 US dollars, with total compensation of around 473 US dollars..
+Reviewers were chosen based on their qualifications, including an average GPA of at least 3.5/4.0 in fundamental math courses, completion or enrollment in the Ordinary Differential Equation course, and meeting English proficiency standards.
+Reviewers evaluated the questions, classifying responses as either numerical answers or "error" if the question was problematic.
+The effectiveness of this review was assessed using Cohen’s Kappa to measure inter-rater reliability, along with individual accuracy rates.
+As shown in Table 3, the pairwise Cohen’s Kappa scores had an average of 0.60, reflecting moderate to substantial agreement among reviewers, with individual scores ranging from 0.50 to 0.71.
+Accuracy rates were moderately high, as detailed in Tables 3.
+For additional details on the review, see Appendix B.
 3.5 Data Statistics
-In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
-The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
-See the word cloud of the combined data in the Appendix L.
+In the section on Ordinary Differential Equations, we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
+The optimization section is divided into two segments: Easy_LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex_LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
+See the category statistics and the word cloud of the combined data in Appendix O.
 4 Evaluation
@@
 4.1.1 Evaluation Details
-Our evaluation protocol is meticulously designed to test the LLMs’ mathematical modeling process.
-For ODE problems, the LLM is prompted to generate Python code from a natural language description.
-The code is executed, and its output is compared with the correct answer to assess accuracy.
-In optimization problems, the LLM is asked to express the model in .lp format, after which the COPT solver is used to find the optimal value for comparison with the correct answer.
+Our evaluation protocol follows the methodology outlined in Section 2.3.
+First, the LLMs were prompted to generate formalized code (in each domain) to represent a mathematical model, denoted as $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$.
+This code was then executed by a solver, and the result, $\widehat{A}$, was compared with the ground truth answer, $A$.
+For ODE problems, the LLM was prompted to generate Python code based on a natural language description.
+The generated code was executed, and the output was compared to the correct solution to assess its accuracy.
+For optimization problems, the LLM was asked to express the model in .lp format, after which an optimization solver COPT (Ge et al., 2023) was employed to find the optimal value for comparison with the correct answer.
+This process is independent of the specific solver used; for instance, MATLAB (Inc., 2022) can be used as an ODE solver, and Gurobi (Gurobi Optimization, LLC, 2024) can serve as an alternative optimization solver.
+Benchmarking Models
+Our evaluation [^2] See the settings in Appendix Q includes both proprietary and open-source LLMs.
+The proprietary LLMs consist of the GPT-4 series (OpenAI, 2023), the Claude series [^3] https://www.anthropic.com/news/claude-3-family, and the Gemini series [^4] https://deepmind.google/technologies/gemini/.
+The open-source LLMs include the DeepSeek series (Shao et al., 2024; DeepSeek-AI, 2024), Llama-3.1 models (Dubey et al., 2024), Qwen-2.5 (Team, 2024), and Mixtral (Jiang et al., 2024).
 Model Output
-In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy).
-In the optimization sections, we prompt LLMs to generate the standard .lp file.
-Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable.
-Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities.
-To better adhere to the format, we conducted tests using 3-shot learning.
-Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix (see Appendix I and Section O).
-Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like ‘ “‘ python “‘ ’ that will affect our testing.
-Mitigation of Formatting Errors
-It is essential to distinguish errors in modeling from those in coding or file formatting [^3] Appendix  H displays the example of formating errors instead of modeling errors..
-Therefore, if the LLM’s output contains syntax errors in python code or .lp code, we need to rectify these issues without altering the underlying logic of the code.
-This step ensures that our evaluation focuses solely on the modeling process.
-To achieve this, we use a language model to fix any syntax error, referring to it as the code modifier.
-In our experiments, we use the original language model and GPT-4-0613 as code modifiers.
-For the sake of reproducibility, we choose the original LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models.
-See the comparison of the result from different code modifiers in Section 6.3, and prompts in Appendix  O.
+We prompted the LLMs to generate Python code that called solvers (such as solve_ivp and odeint in SciPy, and dsolve in NumPy).
+In the optimization sections, the LLMs were asked to generate a standard .lp file.
+Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023), which involved prompting the model to output Python code for invoking the optimization solver, the .lp format more closely aligned with the standard optimization form, making it more readable.
+Optimization solvers like COPT (Ge et al., 2023) could directly read and solve the .lp file.
+To better adhere to this format, we conducted tests using 3-shot learning.
+Detailed information about the number of few-shot samples used in our experiments and the results can be found in Appendix K and Appendix R.
+Additionally, at the start of the testing process, we performed an initial cleaning of the code, removing strings such as ``` python ``` that would have affected our testing.
 4.1.2 Evaluation Metrics
-Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies.
-Let $O$ denote the output from the LLM, and $A$ denote the standard answer.
-The comparison is defined by the following criteria  444 See the evaluation script in Appendix M
-General Comparison Criterion Let $n$ denotes the decimal places of $A$ ($n=0$ if $A$ is integer). The Answer $O$ is correct if
-$\left|\frac{O-A}{A}\right|<10^{-4}\text{ or }|O-A|<\min\{10^{-n},10^{-2}\}$ (1)
-This metric is designed [^5] See the case study in Appendix D to account for the precision errors inherent in solvers and the specified significant figures in questions, ensuring that the LLM’s output is accurately evaluated against the standard answer with allowances for numerical precision and variability in computational responses.
+According to Property 2 and Property 3, both the ground truth answer $A$ and the answer $\widehat{A}$, derived from solving the LLM’s output mathematical model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$, are numerical values with specified significant figures.
+To account for minor discrepancies, we apply the following comparison rule:
+General Comparison Criterion
+: Let $n$ denote the number of decimal places in $A$ ($n=0$ if $A$ is an integer). The answer $\widehat{A}$ is considered correct if:
+$\left|\frac{\widehat{A}-A}{A}\right|<10^{-4}\text{ or }|\widehat{A}-A|<\min\{10^{-n},10^{-2}\}.$
+Figure 4: Examples illustrating the application of the General Comparison Criterion
+This rule accounts for precision errors and ensures an accurate evaluation of $\widehat{A}$ while allowing for slight computational variations.
+Figure 4 demonstrates the application of this criterion with examples, illustrating how discrepancies between the correct and generated answers are assessed.
+For further details, see Appendix P and the case study in Appendix E.
 4.2 Benchmarking Results
-Table 3: Evaluation Results on the Mamo Benchmark.
-The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 6.1.1).
-The "Overall" score represents the weighted average of correct rates, weighted by question count.
-The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier.
-The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
 Models ODE Optimization Overall (%)
-First Second System Easy Complex
-order (%) order (%) (%) LP (%) LP (%)
-GPT-4o 67.86 +0.51 36.36 +0.91 40.00
+First
+Second
+System
+Easy
+Complex
+order (%)
+order (%)
+LP (%)
+LP (%)
+Proprietary Models
+o1-preview 71.43 +0.00 50.00 +0.00 45.00 +0.00 80.21
++0.00
+36.02 +3.79 67.60
++0.66
+GPT-4o 67.86
++0.51
+36.36
++0.91
+40.00
 +0.00
@@
 +8.53
-66.67 +1.73
+66.67
++1.73
 GPT-4-turbo 64.80
@@
 +0.00
-87.88 +0.00 23.22
+87.88
++0.00
+23.22
 +6.64
@@
 +1.81
-45.00
-+2.50
-83.74
+45.00 +2.50 83.74
 +0.31
@@
 +0.00
-30.81 +5.21 66.09
+30.81
++5.21
+66.09
 +1.24
-DeepSeek-v2 39.29
-+1.02
-20.00
-+9.09
-30.00
+Open-source Models
+DeepSeek-v2.5 67.86
 +0.00
-86.50
-+0.16
-3.79
+36.36
 +0.00
-56.49
-+1.08
-DeepSeek-coder-33B 28.57
-+2.04
-11.82
+32.50
 +0.00
-12.50
+86.20
++1.99
+13.33
 +0.00
-75.77
-+1.22
-18.96
-+0.00
-50.29
-+0.99
+65.27
++1.07
 DeepSeek-math-7B-base 3.57
@@
 +0.50
-Llama-3-8b 0.51
+Llama-3.1-8B-instruct 30.61
++1.53
+4.55
++1.81
+10.00
++2.50
+74.39
++2.60
+13.27
++0.00
+48.14
++1.90
+Llama-3.1-70B-instruct 62.24
++0.52
+33.64
++5.45
+32.50
++0.00
+85.74
++0.36
+22.75
++1.42
+64.44
++1.02
+Llama-3.1-405B-instruct 66.84
++1.02
+40.00
++0.00
+30.00
++0.00
+86.20
++0.76
+34.12
++4.74
+67.91 +1.40
+Qwen-2.5-32B-instruct 60.71
 +0.51
-0.00
+38.18
 +0.00
-0.00
+35.00
 +0.00
-55.37
-+0.46
-6.64
+84.05
 +0.00
-31.10
+12.32
++1.42
+61.95
 +0.33
-Llama-3-70b 4.08
-+0.51
-1.82
-+1.82
-2.50
+Qwen-2.5-72B-instruct 60.71
 +0.00
-66.10
-+0.62
-9.00
+37.27
++0.91
+32.50
 +0.00
-38.13
-+0.58
-Qwen-1.5-72b-Chat 5.61
-+6.12
-4.55
-+0.00
-5.00
-+7.50
-65.03
-+6.29
-5.69
-+0.00
-37.55
-+4.64
-Qwen-1.5-110b-Chat 28.57
-+4.08
-6.36
-+1.82
-7.50
-+0.00
-65.18
-+10.89
-13.27
-+2.37
-43.93
-+6.11
+89.42 +0.15 11.43
++1.90
+64.53
++0.50
 Mixtral-8x7B-instruct 23.47
@@
   • 1
-    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125
-Our evaluation  [^6] See the setting in Appendix N includes both closed-source LLMs and open-source LLMs.
-The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^7] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^8] https://deepmind.google/technologies/gemini/.
-The open-source LLMs include DeepSeek series by  DeepSeek-AI (2024),  Guo et al.
-(2024) and  Shao et al.
-(2024); Llama-3 models  [^9] https://github.com/meta-llama/llama3; Qwen-1.5 (Bai et al.) and Mixtral (Jiang et al.).
-Table 3 presents the evaluation results of different language models on the Mamo Benchmark.
-The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (the improvement is indicated by the subscript).
-Table 5 in Appendix A shows results when GPT-4-0613 is selected as the code modifier.
-The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks.
-Specifically, GPT-4o achieves the highest performance in ODE tasks with 68.37% in first-order and 37.27% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 31.28% in complex LP scenarios when using itself as code modifier.
-Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability, achieving a 57.57% success rate across all tasks when enhanced with syntax correction by itself, underscoring its potential despite its open-source nature.
-Moreover, Mixtral-8x22B also exhibits competitive performance relative to DeepSeek-v2.
-After fix syntax error by itself, it achieves a remarkable score of 21.33% on the complex LP task, showcasing its potential in modeling complex optimization problems.
-The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models.
-For instance, GPT-3.5-turbo’s performance in ODE tasks increased from 12.72% to 18.21% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 32.95% to 38.73%.
-Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 32.08% to 35.55% when enhanced with syntax correction.
-6.3 On the Format Errors and Code Modifiers
-Table 4: Execution Success Rates for Raw, Self-Modified, and GPT-4 Modified Methods, denoting scenarios without code modifiers, utilizing the LLM as a code modifier, and employing GPT-4-0613 as the code modifier.
-This table highlights the success rates of file execution (correct format) across these three scenarios, differing from the Table 3 which shows the correctness rate.
+    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125.
+Table 4: Evaluation Results on the Mamo Benchmark.
+The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 4.3).
+The "Overall" score represents the weighted average of correct rates, weighted by question count.
+The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier.
+The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
 Models ODE (%) Optimization (%)
 raw self-modified GPT-4-modified raw self-modified GPT-4-modified
+Proprietary Models
+o1-preview 99.71 100.00 100.00 97.80 99.65 98.73
 GPT-4o 96.82 99.42 99.42 92.93 97.91 97.91
@@
 Gemini-1.5-pro 93.64 98.27 98.84 95.48 98.73 96.87
-DeepSeek-v2 70.81 83.82 96.24 90.27 92.00 94.90
-DeepSeek-coder 88.44 93.64 95.95 92.58 96.29 98.84
+Open-source Models
+DeepSeek-v2.5 97.69 99.13 99.13 90.36 93.83 97.55
 DeepSeek-math-7B-base 28.90 29.77 79.77 55.50 58.40 75.67
 DeepSeek-math-7B-rl 5.20 6.64 80.64 19.93 21.55 55.97
-Llama-3-8B 20.23 26.01 68.21 75.09 76.36 79.14
-Llama-3-70B 24.56 30.92 77.46 77.64 79.49 89.46
-Qwen-1.5-72B-Chat 55.49 89.60 95.38 77.40 92.24 95.37
-Qwen-1.5-110B-Chat 51.45 63.29 95.09 71.26 84.01 90.15
+Llama-3.1-8b-instruct 57.51 72.97 92.20 80.99 88.30 94.55
+Llama-3.1-70b-instruct 92.48 98.27 97.98 92.93 96.99 97.34
+Llama-3.1-405b-instruct 98.84 99.71 99.71 92.58 98.03 95.48
+Qwen-2.5-32B-instruct 97.69 98.84 99.13 93.17 96.29 94.67
+Qwen-2.5-72B-instruct 97.40 97.69 97.98 90.80 94.06 91.15
 Mixtral-8x7B-instruct 50.87 69.08 91.62 84.13 87.83 96.29
 Mixtral-8x22B-instruct 80.92 88.73 96.24 87.72 93.51 96.29
-The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different modification methods.
-The table considers three scenarios: no modification, self-modification, and modification provided by GPT-4-0613.
-The GPT-3.5-turbo model shows the most dramatic improvement when modified by GPT-4, jumping from a 37.28% success rate to 93.93%.
-Interestingly, Claude-3 models (Sonnet and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4-0613, highlighting the compatibility of GPT-4’s modification methods with other models.
-In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
-The modification from GPT-4 is generally superior to both self-modification and no modification.
-This indicates that GPT-4 can serve as a more effective and fair code modifier.
-By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors.
-Table 5 in Appendix A shows the results of selecting GPT-4-0613 as code modifier.
-6.4 Analysis
-Given the results in Section 6.26.3 and Appendix A, several observations emerge from the benchmark results.
-For open-source models, larger models generally perform better, as evidenced by comparisons between Llama-3-70B and Llama-3-8B, Mixtral-8x22B-instruct and Mixtral-8x7B-instruct, with the larger models consistently achieving higher scores.
-From Table 4, it is apparent that some models struggle with adhering to the syntax of the solver, such as GPT-3.5-turbo and certain open-source models.
-Some models like Llama-3-8B may not closely follow the instructions, even worse than the syntax error.These models tend to show significant improvements in both correct rates (as indicated by the subscript in Table 3) and execution success rates after the application of code modifiers, particularly when their initial performance without modifiers is low.
-Notably, in the ODE part, GPT-3.5-turbo achieves a remarkable 28.03% increase in execution success rate by self-modified, accompanied by substantial improvements in ODE performance.
-In Optimization part, Qwen-1.5-72B-Chat follows the same trends.
-When comparing the results from Table 3 and Table 5, it becomes clear that the improvements are even more pronounced for open-source models when utilizing GPT-4 modifications rather than self-modifications.
-See Appendix E for the analysis of typical modeling errors and case study,
+Table 5: Execution Success Rates for Raw, Self-Modified, and GPT-4 Modified Methods, denoting scenarios without code modifiers, utilizing the LLM as a code modifier, and employing GPT-4-0613 as the code modifier.
+This table highlights the success rates of file execution (correct format) across these three scenarios, differing from the Table 4 which shows the correctness rate.
+Table 4 presented the evaluation results (accuracy rate) of different language models on the Mamo Benchmark.
+The performance of these models was assessed both in their original form and when enhanced by the language model (i.e., code modifier) itself to rectify syntax errors (with improvements indicated by the subscript).
+Appendix A provided results when GPT-4-0613 was used as the code modifier, and Appendix F presented the error analysis.
+Mamo is Challenging
+As the results in Table 4 indicate, even state-of-the-art models struggled with complex tasks such as second-order ODEs, systems of ODEs, and Complex_LP.
+This is due to the increasing number of variables and the more intricate relationships among these variables compared to easier tasks like Easy_LP.
+Take-away 1.
+Existing LLMs struggle with complex tasks in mathematical modeling.
+The Performance of o1
+Interestingly, o1-preview, although it excelled in complex tasks such as second-order ODEs and Complex_LP, did not outperform (and was even worse than GPT-3.5-Turbo) in Easy_LP tasks, which consist of high-school-level linear programming modeling problems.
+As shown in Table 6, the execution success rates of o1-preview in each category were nearly 100% across all methods (raw, self-modified, or GPT-4-modified), indicating that this underperformance was not due to limitations in coding ability.
+Take-away 2.
+o1 excelled in complex tasks but underperformed in easy tasks.
+Figure 5: The testing pipeline involving code modifier.
+Open-source Models vs. Proprietary Models
+One of the most notable aspects of this benchmark is the emergence of open-source models as serious competitors to closed-source models.
+The Llama-3.1-405B, in particular, achieved the best overall score across both evaluation categories.
+The Llama-3.1 series, Qwen series, and DeepSeek-v2.5, despite being open-source models, demonstrated remarkable performance, even surpassing highly regarded commercially developed models such as GPT-4 and Claude-3 in categories like second-order ODEs and Optimization.
+Take-away 3.
+Open-source models are competitive in simpler tasks, but there is still a gap compared to closed-source models in more complex tasks.
+Effectiveness of Scaling
+The evaluation demonstrates a correlation between model size and performance across almost all tested categories.
+As indicated by the results of the Llama-3.1, Qwen, and Mixtral series, for models with the same architecture but different parameter scales (e.g., Llama-3.1-7B vs.
+Llama-3.1-405B), larger models tended to outperform their smaller counterparts.
+This observation highlights the effectiveness of scaling in the mathematical reasoning tasks of LLMs.
+However, the effectiveness of scaling may plateau on easier tasks, such as Easy_LP problems.
+Take-away 4.
+Larger LLMs perform better in mathematical modeling.
+4.3 On the Impact of Format Errors
+Mitigation of Formatting Errors
+It is important to differentiate between errors in modeling and those related to coding or file formatting (See Appendix I).
+When the LLM output contains syntax errors in Python or .lp code, these errors must be corrected without altering the logic of the code to ensure that the evaluation focuses solely on modeling.
+To address syntax issues, we employ a code modifier, as the pipeline shown in Figure 5 (More details are discussed in Appendix D).
+In our experiments, we used the original language model and GPT-4-0613 as code modifiers, with the former selected for reproducibility and the latter for its superior coding proficiency and stability.
+A comparison of results from different code modifiers is provided in Section 4.2 and Appendix A, with prompts detailed in Appendix R.
+Setting
+The testing pipeline is shown in Figure 5.
+As mentioned in Section 4.1.1, we selected the original LLM and GPT-4-0613 as code modifiers.
+The data presented in Table 5 provides a comparison of the performance (rate of correct format) of various models using different modification methods: i) no modification, ii) self-modification, and iii) modification provided by GPT-4-0613.
+Observation
+Combining Table 4 and Table 5, the improvement can be attributed to misjudgments caused by the LLM’s coding ability.
+For LLMs with a lower initial correct format rate (such as GPT-3.5-Turbo and Llama-3.1-8B in ODE tasks), the relative improvement is higher.
+This is because, in several cases, their output codes, initially judged as "Error" due to compile errors, actually represent correct mathematical models.
+It is also worth noting that the improvement has little effect on the ranking, indicating that the influence of different coding abilities on our benchmark is minimal.
+Take-away 5.
+Format errors have minimal impact on mathematical modeling.
+5 Related Work
+Recent studies have advanced the use of large language models (LLMs) in mathematical problem-solving by introducing new datasets and solver methodologies.
+For example, Frieder et al.
+(2024) and Yuan et al.
+(2023) developed datasets to test LLMs on various mathematical problems and arithmetic expressions.
+Lewkowycz et al.
+(2022) focused on training models using natural language paired with LaTeX-formatted math from arXiv.
+To address complex computational tasks, Zhang et al.
+(2024) introduced the CARP dataset, while He et al.
+(2024) presented OlympiadBench, a bilingual benchmark for competition-level math reasoning.
+Liu et al.
+(2024) also developed MathBench, structured to assess LLMs’ theoretical knowledge.
+Additionally, Mirzadeh et al.
+(2024) proposed GSM-Symbolic, a benchmark that generates symbolic variants of GSM8K questions.
+Their study also introduced GSM-NoOp, a dataset containing irrelevant but seemingly related information to assess model robustness, demonstrating that while LLMs exhibit sensitivity to numerical variations, they remain resilient to superficial modifications.
+In solver development, efforts have focused on integrating LLMs into novel problem-solving approaches.
+Yang et al.
+(2023) explored LLMs as direct solvers, highlighting their autonomous problem-solving potential.
+He-Yueya et al.
+(2023) introduced a hybrid method that combines LLMs with external symbolic solvers for equation solving, while Pan et al.
+(2023) developed LOGIC-LM, a framework blending LLMs with symbolic solvers to tackle logical problems.
+AhmadiTeshnizi et al.
+(2023) contributed to both dataset development and solver integration by creating NLP4LP, a benchmark for LP and MILP problems, and OptiMUS, an LLM-based agent for optimization problem solving.
+In line with these developments, Lyu et al.
+(2023) introduced Faithful-CoT, a two-stage framework that translates natural language into symbolic reasoning chains (represented as code) before solving the problem using corresponding solvers.
+This approach aligns closely with our benchmark’s objective of assessing the faithful translation of problems into structured mathematical formulations.
 6 Conclusion and Future Directions
-In conclusion, Mamo’s innovative benchmark shifts the evaluation focus of LLMs towards the modeling process, offering a nuanced assessment of their mathematical problem-solving capabilities.
-By isolating modeling from solving through the use of solvers, Mamo provides a comprehensive understanding of LLMs’ strengths and limitations in mathematical modeling.
-This approach not only enhances our insights into the fundamental modeling abilities of LLMs but also guides future research by highlighting the importance of the modeling process in solving mathematical problems.
+Mathematical modeling is a crucial yet challenging aspect of mathematical reasoning.
+Evaluating LLMs’ ability to handle modeling is essential, as their performance will continue to improve over time, making specialized benchmarks in specific areas necessary to accurately assess and push their capabilities.
+Due to the complexity of mathematical models, assessment is difficult.
+To address this, we introduce the Mamo benchmark, which incorporates solvers to validate answers and assess the correctness of the modeling process.
+Mamo demonstrates that LLMs can be effectively evaluated on complex reasoning tasks, where the correct answer may vary but remains equivalent in principle.
+This approach paves the way for further research into LLMs’ reasoning capabilities.
+Limitations
+While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research.
+The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning.
+This opens the door for benchmarks that better distinguish between an LLM’s modeling skills and formalization abilities.
+Future work could focus on assessing LLMs’ mathematical modeling directly, bypassing formalization for a more nuanced evaluation.
+Ethics Statement:
+The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes.
+Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups.
+This includes but is not limited to issues of privacy, security, and the potential for misuse of the model.
+Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.
+Acknowledgment
+This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
 Appendix A GPT-4-0613 as Code Modifier
-[…]
 Models ODE Optimization Overall (%)
@@
 order (%) order (%) (%) LP (%) LP (%)
+Proprietary Models
+o1-preview †† 71.43 50.00 45.00 80.21 37.91 67.49
 GPT-4o †† 68.37 37.27 40.00 87.42 29.38 68.07
@@
 Gemini-1.5-pro †† 65.82 30.91 50.00 85.28 34.12 67.08
-DeepSeek-v2 †† 54.08 30.91 30.00 86.81 3.79 60.05
-DeepSeek-coder †† 30.10 11.82 15.00 78.83 18.96 52.27
+Open-source Models
+DeepSeek-v2.5 †† 67.86 36.36 32.50 86.20 13.33 64.20
 DeepSeek-math-7b-base †† 13.78 9.09 5.00 59.05 9.48 36.72
 DeepSeek-math-7b-rl †† 13.78 7.27 7.50 47.70 1.90 29.20
-Llama-3-8b †† 11.22 0.91 0.00 63.19 8.53 37.47
-Llama-3-70b †† 14.80 7.27 10.00 70.71 9.48 43.18
-Qwen-1.5-72b-Chat †† 16.84 13.01 5.45 78.53 6.64 47.80
-Qwen-1.5-110b-Chat †† 40.82 13.64 20.00 78.83 14.22 53.51
+Llama-3.1-8b-instruct †† 33.67 8.18 15.00 82.06 13.27 53.27
+Llama-3.1-70b-instruct †† 62.76 39.09 32.50 86.81 23.70 65.76
+Llama-3.1-405b-instruct †† 67.86 40.00 30.00 86.66 34.60 68.40
+Qwen-2.5-32b-instruct †† 61.73 38.18 35.00 84.05 13.27 62.28
+Qwen-2.5-72b-instruct †† 60.71 37.27 32.50 89.57 11.43 64.61
 Mixtral-8x7b-instruct †† 32.14 11.82 20.00 75.46 9.95 49.38
 Mixtral-8x22b-instruct †† 55.10 17.27 22.50 85.43 22.27 61.21
+[…]
 Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement.
-For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 32.08% with self-modification to 43.93% when GPT-4 is used as the code modifier.
+For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 36.41% with self-modification to 48.80% when GPT-4 is used as the code modifier.
 This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
@@
 We have also conducted the cross-review process. See the details in Appendix B.1 and the results on Appendix B.1.1.
-In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions.
+In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and six had misleading descriptions.
 All invalid entries were filtered out, and the remaining issues were corrected manually.
-Among them, 8 of them were corrected and others (36 questions) were filtered out.
+Among them, eight of them were corrected and others (36 questions) were filtered out.
 In complex LP part, for the original 211 questions, 40 were found formulated by wrong mathematical models.
@@
 Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis.
-Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^10] The hourly wage pay is 100 RMB/h, the total amount spent on the participant compensation is 2400 RMB, approximately 331 dollars.
+Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^5] The hourly wage is approximately 20 US dollars (equivalent in local currency), with a total participant compensation of around 473 US dollars (equivalent in local currency)..
 The responses could either be numeric or labeled as “error” if a question was deemed problematic.
@@
 These results provide insights into the overall effectiveness and areas of improvement for the cross-review process, particularly in terms of aligning understanding and interpretation of the evaluation criteria among reviewers.
-Appendix C Annotator Qualifications
+Appendix C Collector Qualifications
 The benchmarking process benefits from the expertise of reviewers with robust backgrounds in mathematics, underscored by their academic accomplishments.
@@
 These criteria ensure their ability to read questions in English and solve ODE problems.
-Appendix D Metric
+Appendix D Understand Code Modifier
+Figure 5 shows the evaluation process involving the code modifier.
+The introduction of code modifiers is to maintain the fairness of the evaluation: we try to reduce conditions when the LLM outputs the code with a correct mathematical model, but the code ends with a compile error due to the limit of coding ability.
+We concider two factors when applying code modifier:
+  1. (a)
+    How often does a format error occur?
+  2. (b)
+    How many passes on average are needed?
+The factor (a) is exactly the Table 5.
+As for (b), the number of passes required to fix format errors depends on the code modifier chosen.
+To study the factor (b), To study factor (b), we applied GPT-4-0613 as a code modifier on the chosen format error files and iteratively called it until the code was executed successfully.
+Table 7 displays the distribution of the number of pass the code modifier was called until the errors were fixed.
+The chosen files were generated by GPT-4o and DeepSeek-v2.5, representing a close-source and an open-source model, respectively.
+1 2 3 4 $\geq$5 all
+LP 74 35 10 6 11 136
+ODE 88 3 1 1 9 112
+Table 7: Number of pass to apply code modifier
+On average, 1.68 passes were needed to correct format errors when using GPT-4-0613.
+We found that a single pass was sufficient to correct the majority of format errors.
+To balance efficiency and accuracy, we only applied a single pass for the code modifier throughout our experiments.
+Appendix E Metric
 The comparison process is as the following:
-$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (1)} or }|O-A|<1\times min\{10^{-n},10^{-2}\}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (2)},}\\ 0&\text{otherwise.}\end{cases}$
+[…]
 Our consideration: (1) The difference is much less than the Answers, so can be omitted.
 (2) The only difference lies in the last few digits of the required significant figure of the answer.
-By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|O-A|<\xi\times min\{10^{-n},10^{-2}\}$.
+By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|\widehat{A}-A|<\xi\times min\{10^{-n},10^{-2}\}$.
 We have explored the relationship between various thresholds and the accuracy of GPT-4o.
@@
 Complex LP 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75
-Figure 2: Accuracy of different thresholds in different categories
+Figure 6: Accuracy of different threshold in different category
 Figure 6 plots the results. In our metric, we choose strict $\xi=1$.
-Appendix E Error Analysis
+Appendix F Error Analysis
 In this section, we will analyze some typical errors.
-E.1 Deviation from Instruction
+F.1 Deviation from Instruction
 It is common for some models that show a low capability of few-shot learning, especially for small models.
 For example, in a case Llama-3-8B output ‘I wish you good luck’ when prompting with some few-shot demonstrations in mathematical modeling.
-E.2 Syntax Error
+F.2 Syntax Error
 In ODE parts, the syntax errors is obviously the syntax errors in the reponsed python code.
 In Optimization part, the .lp file require more strict syntax than the universal optimization formulation.
-The Figure 4 and 4 show the typical syntax errors in .lp format.
+The Figure 8 and 8 show the typical syntax errors in .lp format.
 Figure 7: The syntax error is caused by the constraint, in .lp format, the variables should be placed on the left hand side of the inequality
 Figure 8: Same type of syntax error as Figure 8
-E.3 Modeling Error
+F.3 Modeling Error
 One of the most common errors in modeling is the lack of recognition of decision variable types.
@@
 Figure 9: In this TSP problem, the decision variable should be binary, instead of "General", which means “integer”
-Appendix F Scope of Mamo
-Table 7: The taxonomy of mathematical modeling. In the Mamo benchmark, we only include the categories where sophisticated solvers are available.
+Appendix G Scope of Mamo
 Category In Mamo
@@
 Probabilistic Modeling
+Table 9: The taxonomy of mathematical modeling. In Mamo benchmark, we only include the categories where sophisticated Solvers is available.
 The scope of Mamo is chosen under the consideration in Section 3.1. As Table 9 shows, Mamo focus on the ODE and Optimization areas.
-Appendix G Examples of Different ${\widehat{M}}$
+Appendix H Examples of Different ${\widehat{M}}$
 Different models will typically lead to different final answers after solving by the solvers. For example, in ODE part:
@@
 The minor difference in the mathematical models will result in different final values.
-Appendix H Examples of Syntax Errors
+Appendix I Examples of Syntax Errors
 Some outputs of LLM contain the correct model but not follow the syntax of solvers. For examples, the standard .lp file for
@@
 However, if we replace the constraint $c1$ with x <= 20 - y, the code will be reported as an error, even if the two constraints are equivalent.
-To specifically test the modelling ability of the LLMs, we apply code modifier to fix the syntax errors.
-Appendix I Few-shot Experiment
+To specifically test the modeling ability of the LLMs, we apply code modifier to fix the syntax errors.
+Appendix J Example of Mathematical Model Correction
+In Figure 10, we display an example of correction of the mathematical model in Complex_LP.
+In the optimization model of the max flow problem, variables were initially set in the wrong signs in two constraints($C_{1},C_{3}$).
+The corrected constraints($C_{1}^{\prime},C_{3}^{\prime}$) are also displayed.
+[…]
+Figure 10: The example of model correction
+Appendix K Few-shot Experiment
 To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance.
@@
 The accuracy of GPT-4-0613 was 66.56% in Easy LP and 14.69% in Complex LP, resulting in an overall LP accuracy of 53.65%, consistent with the results obtained from sampling.
-Appendix J Testing Process
-We use 3-shot learning in testing; see examples in Appendix O. The examples of testing process in ODE and optimization is shown in Figure 8 and Figure 9.
+Appendix L Testing Process
+We use 3-shot learning in testing; see examples in Appendix R. The examples of testing process in ODE and optimization is shown in Figure 13 and Figure 14.
 Figure 13: Example of testing process in optimization
 Figure 14: Example of testing process in ODE
-Appendix K Question Formulation
-Figure  10 shows the example of synthesis process in ODE.
+Appendix M Probability of Getting Correct $\widehat{A}$ from Wrong ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$
+There are chances for the wrong mathematical model $\widehat{M}$ to lead to correct answer $\widehat{A}$.
+However, the probability is low due to the wide spread of the continuous answer space and lack of prior knowledge.
+We examined the distribution of all the answers in Mamo dataset.
+The probability of guessing an ODE answer within a 0.01 interval is about 2.9%, and for LP, it’s around 7.3%.
+These low probabilities highlight the difficulty of correctly guessing the answer without proper modeling.
+Appendix N Question Formulation
+Figure  15 shows the example of synthesis process in ODE.
+The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity.
+This is followed by the computation of answers to establish a ground truth.
+Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems.
 Figure 15: Example of synthesis process in ODE
-Appendix L Word Cloud
-See the word cloud of the combined data in Figure 11
+Appendix O Data Description and Word Cloud
+We present a comparative table showcasing the number of scenarios across various categories, from manufacturing to natural science, in Table 10.
+We also make the word cloud of the combined data in Figure 16
+Category Manufacturing Environment Business Daily Life Natural Science Others
+LP 75 114 265 258 44 77
+ODE 85 74 54 22 109 2
+Table 10: Comparison of ODE and LP dataset of different categories
 Figure 16: The word cloud of the data after filtering the requirement words such as ’significant figures’, ’rounded’
-Appendix M Evaluation script
+Appendix P Evaluation script
+The evaluation script is as following:
 def comp(output, standard_answer):
@@
         return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
-Appendix N Settings
+Appendix Q Settings
 We have the following settings:
-  • We testing the following models by calling API: All the closed source models, DeepSeek-V2 and DeepSeek-coder , with temperature=0.8
+  • We tested the following models by calling API: All the closed source models except o1-preview, DeepSeek-V2.5, Llama-3.1 series, Qwen-2.5 series with temperature=0.8.
+o1-preview was called under default temperature since the change of temperature is not allowed.
   • Other models was deployed and inference using VLLM (Kwon et al., 2023) on machine in slurm cluster with 8 * A100, with temperature 0.8 and max_length 4096 tokens.
-Appendix O Prompts
+Appendix R Prompts
+The following figures show the prompts we use in the benchmarking.
+Figure 17,18 display the prompt we use in the evaluation process, while Figure 19,20 list the prompts we use for calling code modifiers.
+As shown in Figure 19,20, the code modifier is prompted to adjust the code based solely on the code and error information during execution, without any information about the problem itself.
+This ensures that the code modifier does not alter the underlying mathematical model based on problem details.
 [Instruction] Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems.
@@
 - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables.
-Here comes the examples: [Example_1] (the input) A manufacturing company produces two types of products: $X$ and $Y$.
+Here comes the examples: [Example_1:] (the input) A manufacturing company produces two types of products: $X$ and $Y$.
 The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$.
@@
 Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution.
-Here comes the [examples] : […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements.
+Here comes the examples: [Examples] […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements.
 And Please directly generate the code without any explaintion(except the comments in the code).
@@
 Figure 20: The prompt for fixing syntax error in LP. [lp_code] refer to the lp code which contains syntax error
-Appendix P Limitations
-While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research.
-The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning.
-This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities.
-Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
-Appendix Q Ethics Statement:
-The benchmark is designed with a stringent ethical framework to ensure that all problems are socially responsible and do not perpetuate or encourage harmful biases or stereotypes.
-Questions are meticulously reviewed to avoid any content that could lead to the dissemination of misinformation, support unethical practices, or cause harm to individuals or groups.
-This includes but is not limited to issues of privacy, security, and the potential for misuse of the model.
-Furthermore, the benchmark abstains from any problem that could indirectly endorse unethical behaviors or decisions in real-world scenarios, upholding the highest standards of academic integrity and social responsibility.
+Appendix S Licensing Information
+This content(incolding benchmark) is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
+Appendix T Explanation of Criteria
+In this section, we explain the reasoning behind the validation of each proposition shown in Figure 3, based on the properties outlined in our benchmark dataset.
+T.1 Property 1: Final-State Approach Solvability
+Property 1 states that problems should be solvable directly, using a ’final-state approach’.
+This means that once the model is formulated, the solver should be able to compute the final answer without requiring additional transformations.
+We assess two propositions based on this property:
+  • Proposition 1: $A=y(t_{0})$ is valid as it asks for the value of the solution function $y(t)$ at a specific time $t_{0}$, which is a final-state solution. A solver can compute this directly.
+  • Proposition 1: $A=a\times y(t_{0})+b$ is invalid because it requires an additional transformation (i.e., multiplying by $a$ and adding $b$), which does not adhere to the final-state solvability approach.
+T.2 Property 2: Unified and Numerical Answers
+Property 2 requires the answer to be a unified and numerical value, particularly for ODE problems where the function’s value at a specific time is sought.
+For optimization problems, the optimal value should be clear.
+  • Proposition 2: $A=\max(x+y)$ for $x,y\leq 0$ = 0.0 is valid as it yields a clear numerical result, specifically the maximum of $x+y$ within the defined constraints.
+  • Proposition 2: $A=\sqrt{3}$ is invalid because while $\sqrt{3}$ represents a number, it is not expressed as a numerical answer.
+In practice, this could lead to ambiguities if different solvers provide decimal approximations (e.g., $1.732$) instead of the exact expression.
+T.3 Property 3: Significant Figures and Precision
+Property 3 mandates that each question specifies the required precision or number of significant figures for the answer.
+  • Proposition 3: $A=6.66$ is valid because it presents the number to three significant figures, which is explicitly required.
+  • Proposition 3: $A=6.7$ is invalid as it is given only to two significant figures, which does not meet the required precision.
+T.4 Property 4: Real-World Problem Context
+Finally, Property 4 introduces the need for real-world context in problem framing, without explicitly stating the mathematical model.
+This requires the solver to abstract the appropriate mathematical model from the problem description.
+  • Proposition 4: "In a company…" aligns with this property, as the problem is framed as a real-world scenario.
+The mathematical model is not explicitly given, challenging the solver to formulate it based on the provided context.